### v2

# On Root Capacity, Intersection Indicium, Minimal Generating Sets of Galois Closure & Compositum Feasible Triplets

Abstract.
We carry forward the work started by the author and Bhagwat in [2] and develop the Theory of root clusters further in this article and also apply similar methods to resolve certain problems in related areas. We establish the Inverse root capacity problem for number fields which is a generalization of Inverse cluster size problem for number fields proved in [2]. We introduce the concept of intersection indicium as a generalization of the concept of ascending index introduced in [2] and prove some of its properties. We then establish the Inverse intersection indicium problem for number fields (excluding certain cases) which is a generalization of Inverse ascending index problem for number fields proved in [2]. We give a field theoretic formulation for the concept of minimal generating sets of splitting fields which was introduced by the author and Vanchinathan in [10] and generalize a result in [10] for number fields and also establish the existence of field extensions over number fields for given degree and given cardinality of minimal generating set of Galois closure dividing the degree. We generalize a result in [6] by establishing that a certain family of triplets is compositum feasible over any number field and we also list all the irreducible triplets in this family. We also prove a partial case of a conjecture in [6]. In the concluding section of this article, we improve on the inverse problems proved in [2] and this article by proving that there exist arbitrarily large finite families of pairwise non-isomorphic extensions having additional properties that satisfy the given conditions.
2020 Mathematics Subject Classification:
11R04, 11R21, 11R32, 12F05, 12F10, 20B35, 20F16
1. Introduction
The Theory of root clusters was substantially developed by the author and Bhagwat in their work in [2] which built on previous work by Perlis in [15] and Krithika and Vanchinathan in [11]. This article is yet another contribution in enriching the theory further by means of generalizations and we also apply similar methods to resolve certain problems in related areas.
Let $K$ be a perfect field. We fix an algebraic closure $\bar{K}$ once and for all and work with finite extensions of $K$ contained in $\bar{K}$. Let $L/K$ be a degree $n$ extension and $\tilde{L}$ be its Galois closure inside $\bar{K}$. Let $G={{\rm Gal}}(\tilde{L}/K)$ and $H={{\rm Gal}}(\tilde{L}/L)$. We have the notion of cluster size of $L/K$, $r_{K}(L)$ which is $N_{G}(H):H]$ (See Section 2.1 in [2] for basic properties of cluster size of field extension). From Section 3.2 in [2], number of clusters of $L/K$, $s_{K}(L)$ is $G:N_{G}(H)]$ which is also the number of distinct fields inside $\bar{K}$ isomorphic to $L$ over $K$.
In Section 2, we prove some interesting properties of unique intermediate extensions for given extensions. This notion was introduced in Section 7 of [2]. The concepts of strong cluster magnification and root capacity were introduced by the author and Bhagwat in Sections 4 and 6 of [2] respectively. In Section 3, we establish the Root Capacity Magnification Theorem, Theorem 3.6 which is a generalization of Cluster Magnification Theorem proved in [11]. The concept of cluster towers was introduced in [11]. We give a field theoretic formulation for cluster towers and prove Proposition 3.8 about strong cluster magnification and cluster towers.
In Section 5, we establish Theorem 5.2 as a generalization of Inverse cluster size problem for number fields (Theorem 3.1.1 in [2]) which is as follows. For notations see Sections 3 and 5.
Theorem 1.1.
(Inverse Root Capacity Problem for Number Fields) Let $K$ be a number field. Given $(n,r,\rho)$ where $n>2$ and $r|n$ and $r|\rho$ and $\rho\neq n-1$. There exist extensions $L/K$ and $M/K$ such that $[L:K]=n$ and $r_{K}(L)=r$ and $\rho_{K}(M,L)=\rho$. For $\rho\neq 0$, we get $M/K$ as an extension of $L/K$ contained in $\tilde{L}$.
The concept of ascending index of an extension was introduced by the author and Bhagwat in Section 7 of [2]. Just like we introduced root capacity as a generalization of cluster size in [2], in Section 4 of this article we introduce the concept of intersection indicium as a generalization of ascending index. We prove some properties of intersection indicium and then establish the Intersection Indicium Magnification Theorem, Theorem 4.7 which is a generalization of Ascending Index Magnification Theorem proved in [2]. In Proposition 4.8, we compute root capacity and intersection indicium for an interesting example.
In Section 6, we establish Theorem 6.2 as a generalization of Inverse ascending index problem for number fields (Theorem 9.0.5 in [2]) which is as follows. For notations see Sections 4 and 6.
Theorem 1.2.
(Inverse Intersection Indicium Problem for Number Fields) Let $K$ be a number field. Given $(n,t,\tau)$ where $n>2$ and $t\ |\ \tau\ |\ n$ (except for the cases listed in Remark 6.6.1). There exist extensions $L/K$ and $M/K$ such that $[L:K]=n$ and $t_{K}(L)=t$ and $\tau_{K}(M,L)=\tau$. We get $M/K$ as an extension of $L/K$ contained in $\tilde{L}$.
The notion of minimal generating sets of the splitting field of a polynomial was introduced by the author and Vanchinathan in Section 2 in [10]. In Section 7, we give a field theoretic formulation for minimal generating sets and prove Proposition 7.4 about strong cluster magnification and minimal generating sets. We prove Theorem 7.6 as a generalization of Theorem 2.11 in [10] for number fields. We then go on to establish Theorem 7.7 (by applying similar methods as in earlier sections) which is as follows. For notations see Section 7.
Theorem 1.3.
Let $K$ be a number field. Given positive integers $n>2$ and $s|n$ with $s<n$. There exists an $L/K$ of degree $n$ for which the Galois closure has a minimal generating set of cardinality $s$.
Furthermore that $L/K$ satisfies $s_{K}(L)=s$. Hence there is a unique minimal generating set for the Galois closure of $L/K$ which is thus, also a minimum minimal generating set.
Compositum feasible triplets as a notion was introduced by Drungilas et al. in [6] along with notions of sum feasible and product feasible triplets, with $\mathbb{Q}$ as the base field. In our discussion in Section 8, we take base field $K$ to be any perfect field unless specified otherwise. We establish Theorem 8.7 (by applying similar methods as in earlier sections) as a generalization of Theorem 7 in [6] for number fields, which is as follows. For notations refer to Section 8.
Theorem 1.4.
Let $K$ be a number field. Given positive integers $a,b$ and t|gcd(a,b)t\ |\ gcd(a,b) we have that $(a,b,lcm(a,b)\cdot t)$ is a compositum feasible triplet over $K$. Furthermore we can choose the field extensions such that compositum has Galois closure over $K$ having a solvable Galois group.
We also have that $(a,b,lcm(a,b)\cdot t)$ is a sum feasible and product feasible triplet over $K$.
Then we prove Proposition 8.10 which is a partial case of Conjecture 4 in [6] where it is conjectured that the set of compositum feasible triplets over $K$ is a commutative multiplicative monoid. The notion of a compositum feasible triplet being irreducible was introduced in [13]. In Proposition 8.14, we list all irreducible triplets in the family of compositum feasible triplets covered in Theorem 8.7. Propositions 8.12 and 8.15 cover some other interesting cases of compositum feasible triplets and sum feasible triplets.
In Section 9, we improve on the inverse problems proved in [2] and this article by proving that there exist arbitrarily large finite families of pairwise non-isomorphic extensions having additional properties that satisfy the given conditions. The following is Theorem 9.2 which is an improvement on Inverse cluster size problem for number fields Theorem 3.1.1 in [2].
Theorem 1.5.
Let $K$ be a number field. Let $n>2$ and $r|n$. Then we get arbitrarily large finite families of extensions $L/K$ inside $\bar{K}$ which are pairwise non-isomorphic over $K$ and are pairwise linearly disjoint over $K$ and each having degree $n$ with cluster size $r_{K}(L)=r$.
We obtain similar improvements for Theorem 9.0.5 in [2] and Theorems 5.2, 7.6, 7.7 and 8.7 which are Theorems 9.3, 9.4, 9.5, 9.6 and 9.7 respectively.
2. Some Remarks on Unique Intermediate Extensions
Let $N/K$ be the unique intermediate extension of $L/K$ such that $L/N$ is Galois with maximum possible degree as in Section 7.1 of [2].
Proposition 2.1.
Consider $L/P/K$.
  1. (1)
    Then $r_{K}(L)/r_{P}(L)=[N_{G}(H):N_{G_{0}}(H)]$ where $G_{0}=Gal(\tilde{L}/P)$. In particular $r_{P}(L)|r_{K}(L)$.
  2. (2)
    Let unique intermediate extension for $L/P$ be $N_{1}/P$. Then $NP=N_{1}$.
  3. (3)
    $s_{K}(L)\ |\ (s_{P}(L)[P:K])$.
Proof.
We will prove (1) and (2).
  1. (1)
    Now $\tilde{L}/P$ is Galois with Galois group $G_{0}$. From the first proposition in [14], $r_{P}(L)=|Aut(L/P)|=[N_{G_{0}}(H):H]$. Thus $r_{K}(L)/r_{P}(L)=[N_{G}(H):H]/[N_{G_{0}}(H):H]=[N_{G}(H):N_{G_{0}}(H)]$. Since $N_{G_{0}}(H)\subset N_{G}(H)$, we have $r_{P}(L)|r_{K}(L)$.
  2. (2)
    By Theorem 7.1.1 in [2], $N=\tilde{L}^{N_{G}(H)}$ and $N_{1}=\tilde{L}^{N_{G_{0}}(H)}$. Also $N_{G}(H)\cap G_{0}=N_{G_{0}}(H)$. Thus $NP=N_{1}$.
Let $\sigma_{i}$ be coset representatives of $N_{G}(H)$ in $G$ with $\sigma_{1}=1$. Let $H_{i}=\sigma_{i}H\sigma_{i}^{-1}$. Then $L_{i}=\tilde{L}^{H_{i}}$ are the $s_{K}(L)$ many distinct fields isomorphic to $L$ over $K$.
Let $F/K$ be the unique intermediate extension of $L/K$ which is Galois with maximum possible degree as in Section 7.2 of [2].
Proposition 2.2.
  1. (1)
    $F=\cap_{i=1}^{s}L_{i}$ where $s=s_{K}(L)$.
  2. (2)
    Let $L=K(\alpha)$ for a primitive element $\alpha\in\bar{K}$ with minimal polynomial $f$ over $K$. A cluster of $f$ is defined as the subset of roots of $f$ belonging to the field generated by a single root of $f$ over $K$ and clusters form equivalence classes for the obvious equivalence relation among the roots of $f$ (For details of this notion, see [11] and [15]). Let $\{\beta_{i}\}_{=1}^{s}$ be a complete set of representatives of root clusters of $f$. Then $F=\cap_{i=1}^{s}K(\beta_{i})$.
  3. (3)
    $F/K$ is the unique intermediate extension of each $L_{i}/K$ which is Galois with maximum possible degree.
Proof.
  1. (1)
    We have from Theorem 7.2.1 in [2], that $F=\tilde{L}^{H^{G}}$ where $H^{G}$ is the normal closure of $H$ in $G$, i.e. the intersection of all normal subgroups of $G$ that contain $H$. We observe that $H^{G}$ is the subgroup of $G$ generated by all $H_{i}$’s for $1\leq i\leq s$. Hence by Galois correspondence, $F=\cap_{i=1}^{s}L_{i}$.
  2. (2)
    Now number of clusters of $f$ is $s_{K}(L)=s$. By a suitable reordering, we have that each $L_{i}=K(\beta_{i})$.
  3. (3)
    Follows from part (1).
Remark 2.2.1.
Proposition 2.2 helps us to give alternate proofs for certain cases in Section 7.3 in [2] which is encapsulated as the following example.
Example 2.3.
  1. (1)
    Consider $K=\mathbb{Q}$ and let $n>2$. Fix $\zeta$ to be a primitive $n$-th root of unity in $\bar{\mathbb{Q}}$. Let $c$ be a positive rational number such that $f=x^{n}-c$ is an irreducible polynomial over $\mathbb{Q}$. Let $a=c^{1/n}$ be the positive real root of $f$. Assume $n$ to be either odd; or even with $\sqrt{c}\not\in\mathbb{Q}(\zeta)$ (Similar to conditions in Example 5.1.3 in [2]). Then the unique $F/\mathbb{Q}$ for $\mathbb{Q}(a)/\mathbb{Q}$ is $\mathbb{Q}$ for $n$ odd & $\mathbb{Q}(a^{n/2})$ for $n$ even.
    By Proposition 1 in [8] and Theorem A in [8], $n$ is odd or, $n$ is even with $\sqrt{c}\not\in\mathbb{Q}(\zeta)$ if and only if $\mathbb{Q}(a)\cap\mathbb{Q}(\zeta)=\mathbb{Q}$. Thus $\mathbb{Q}(\zeta)(a):\mathbb{Q}(\zeta)]=n$. Hence we have the set $\{a^{i}\}_{i=0}^{n-1}$ to be linearly independent over $\mathbb{Q}(\zeta)$. Let $\gamma\in\mathbb{Q}(a)\cap\mathbb{Q}(a\zeta)$.
    Then $\gamma=a_{0}+a_{1}a+\dots+a_{n-1}a^{n-1}=b_{0}+b_{1}(a\zeta)+\dots+b_{n-1}(a\zeta)^{n-1}$ for $a_{i},b_{i}\in\mathbb{Q}$ for all $0\leq i\leq n-1$. Thus $a_{0}+a_{1}a+\dots+a_{n-1}a^{n-1}=b_{0}+(b_{1}\zeta)a+\dots+(b_{n-1}\zeta^{n-1})a^{n-1}$. Hence for all $0\leq i\leq n-1$, we have $a_{i}=b_{i}\zeta^{i}$. If $n$ is odd, $a_{i}=0$ for all $i\neq 0$ and if $n$ is even, $a_{i}=0$ for all $i\neq 0,n/2$. Also for $n$ even, $a^{n/2}\in\mathbb{Q}(a\zeta^{i})$ for all $0\leq i\leq n-1$. Hence by Proposition 2.2, for n odd, $F=\mathbb{Q}$ and for n even $F=\mathbb{Q}(a^{n/2})$.
  2. (2)
    Let $f$ over $K$ be irreducible of degree $n>2$ with Galois group ${\mathfrak{S}}_{n}$ with roots $\alpha_{i}\in\bar{K}$ for $1\leq i\leq n$. Let $L=K(\alpha_{1})$. Then the unique $F/K$ for $L/K$ is $K$.
    Since the Galois group is ${\mathfrak{S}}_{n}$, we have $K(\alpha_{1},\alpha_{2}):K(\alpha_{1})]=n-1$. Hence $K(\alpha_{1})$ and $K(\alpha_{2})$ are distinct fields. Let $[F:K]=t$. By Proposition 2.2, $F\subset K(\alpha_{2})$. We have $K(\alpha_{1}):F]=[K(\alpha_{2}):F]=n/t$. Since $K(\alpha_{1},\alpha_{2}):K(\alpha_{1})]\leq[K(\alpha_{2}):F]$. Thus $n-1\leq n/t$. Which holds only if $t=1$ i.e. $F=K$.
For an extension $L/K$, we have the notion of ascending index of $L/K$, $t_{K}(L)$ which is $G:H^{G}]$ and the quantity $u_{K}(L)$ which is $H^{G}:H]$ (See Section 7.2 and Section 9 in [2] for basic properties of ascending index of field extension).
Proposition 2.4.
Consider $L/P/K$.
  1. (1)
    Then $t_{K}(L)/t_{K}(P)=[G_{0}^{G}:H^{G}]$ where $G_{0}=Gal(\tilde{L}/P)$. In particular $t_{K}(P)|t_{K}(L)$.
  2. (2)
    Let unique intermediate extension for $L/P$ be $F_{1}/P$. Then $F\subset F_{1}$. Thus $t_{P}(L)[P:K]=t_{K}(L)[H^{G}:H^{G_{0}}]$. In particular $t_{K}(L)\ |\ (t_{P}(L)[P:K])$.
  3. (3)
    $u_{P}(L)|u_{K}(L)$ and $u_{K}(L)\ |\ (u_{K}(P)[L:P])$.
Proof.
We will prove (2). Since $H^{G}\cap G_{0}\unlhd G_{0}$. Hence $H^{G_{0}}\subset H^{G}\cap G_{0}\subset H^{G}$. From Theorem 7.2.1 in [2], $F=\tilde{L}^{H^{G}}$ and $F_{1}=\tilde{L}^{H^{G_{0}}}$. Thus $F\subset F_{1}$. Now $t_{P}(L)=[G_{0}:H^{G_{0}}]=[F_{1}:P]$. Since $t_{K}(L)=[G:H^{G}]=[F:K]$ and $F:K][F_{1}:F]=[F_{1}:K]=[F_{1}:P][P:K]$, we are done. ∎
3. Root Capacity and Cluster Towers
We begin with the following definition.
Definition 3.1 (Definition 4.1.1 [2]).
A finite extension $M/K$ is said to be obtained by strong cluster magnification from a subextension $L/K$ if we have the following:
  1. (1)
    $[L:K]=n>2,$
  2. (2)
    there exists a finite Galois extension $F/K$ such that the Galois closure $\tilde{L}$ of $L/K$ in $\bar{K}$ and $F$ are linearly disjoint over $K$ i.e. $\tilde{L}\cap F=K$.
  3. (3)
    $LF=M$.
The number $[F:K]$ is called the magnification factor and denoted by $d$.
The following example negatively answers the Problem 10.2.6 on strong cluster magnification in Chapter 10 of PhD Thesis [9] of the author.
Example 3.2.
Consider the case in Example 2.3 (1) with $n\equiv 2\ (\text{mod}\ 4)$. From Example 7.3.3 in [2], we have that $\mathbb{Q}(a)/\mathbb{Q}$ is obtained by nontrivial strong cluster magnification from $\mathbb{Q}(a^{2})/\mathbb{Q}$ through $\mathbb{Q}(a^{n/2})/\mathbb{Q}$. We also have that both the extensions $\mathbb{Q}(a^{2})/\mathbb{Q}$ and $\mathbb{Q}(a)/\mathbb{Q}(a^{n/2})$ have cluster size 1.
Hence we conclude that if $M/K$ is obtained by nontrivial strong cluster magnification from some $L/K$ and $K\subset M^{\prime}\subset M$ and $K\subset K^{\prime}\subset M$. Then it is not necessary that $M^{\prime}/K$ or $M/K^{\prime}$ are obtained by nontrivial strong cluster magnification from some subextensions.
The following is a reformulation (in terms of strong cluster magnification property) of a result proved for polynomials in Section 3.1 of [11] and reformulated for field extensions in Section 4 of [11] which is referred to as the Cluster Magnification theorem.
Theorem 3.3.
Let $M/K$ be obtained by strong cluster magnification from $L/K$ with magnification factor $d$. Then [M:K]=d[L:K][M:K]=d\ [L:K] and $r_{K}(M)=d\ r_{K}(L)$.
Definition 3.4 (Definition 6.2.1 [2]).
Let $L/K$ be an extension. By primitive element theorem $L=K(\alpha)$ for some $\alpha\in\bar{K}$. Let $f$ be minimal polynomial of $\alpha$ over $K$.
For an extension $M/K$, root capacity of $M$ with respect to $L$ (with base field $K$ fixed) $\rho_{K}(M,L)$ is the number of roots of $f$ that are contained in $M$. (This is well defined by Proposition 6.2.2 in [2]).
Equivalently by Proposition 6.2.6 (1) in [2], $\rho_{K}(M,L)=a\ r_{K}(L)$ where $a$ is number of distinct fields inside $M$ isomorphic to $L$ over $K$.
Remark 3.4.1.
Suppose $M/K$ and $M^{\prime}/K$ are isomorphic over $K$ and $L/K$ and $L^{\prime}/K$ are isomorphic over $K$. Then $\rho_{K}(M,L)=\rho_{K}(M^{\prime},L^{\prime})$.
Example 3.5.
Adding to Example 6.2.7 in [2]. Consider the case in Example 2.3 (2). Let $L_{k}=K(\alpha_{1},\dots,\alpha_{k})$ and let $L=L_{1}$ and $L_{0}=K$. So Galois closure of $L/K$ is $\tilde{L}=L_{n-1}$.
Now $L_{k+1}=L_{k}(\alpha_{k+1})$. One can verify that the minimal polynomial of $\alpha_{k+1}$ over $L_{k}$ has degree nkn-k and has the roots $\alpha_{k+1},\alpha_{k+2},\dots,\alpha_{n}$. Also $\alpha_{i}\not\in L_{k+1}$ for $i>k+1$ and $k\leq n-3$. Thus for $0\leq k\leq n-2$, Galois closure of $L_{k+1}/L_{k}$ is $L_{n-1}/L_{k}$ with Galois group $\mathfrak{S}_{n-k}$.
We have for $0\leq k\leq n-3$ that $L_{k+j}:L_{k}]=\ ^{n-k}P_{j}$ and $\rho_{L_{k}}(L_{k+j},L_{k+1})=j$ where
$1\leq j\leq n-2-k$. Also by Theorem 3 in [11], we have $r_{L_{k}}(L_{k+j})=j!$. Thus by proof of Theorem 3.2.4 in [2], we have $\rho_{L_{k}}(L_{k+j},L_{k+l})=\ ^{j}C_{l}\ r_{L_{k}}(L_{k+l})=\ ^{j}C_{l}\ l!=\ ^{j}P_{l}$ where $1\leq l\leq j$.
We have $r_{L_{k}}(L_{k+j})=j!=l!\ (j-l)!\ ^{j}C_{l}=r_{L_{k}}(L_{k+l})r_{L_{k+l}}(L_{k+j})\ ^{j}C_{l}$. Also for $1\leq m<l$, $\rho_{L_{k}}(L_{k+j},L_{k+l})=\ ^{j}P_{m}\ ^{j-m}P_{l-m}=\rho_{L_{k}}(L_{k+j},L_{k+m})\rho_{L_{k+m}}(L_{k+j},L_{k+l})$.
Remark 3.5.1.
Consider $L/K$ and $M/M^{\prime}/K$. Then $\rho_{K}(M^{\prime},L)\ |\ \rho_{K}(M,L)$ is not true in general. For example consider the case in Example 3.5. Now $\rho_{K}(L_{j},L_{1})=j$ and $\rho_{K}(L_{j+1},L_{1})=j+1$. Now $j\nmid(j+1)$ for $j\geq 2$.
Remark 3.5.2.
Consider $M/L/L^{\prime}/K$. Then the statements $r_{K}(L^{\prime})r_{L^{\prime}}(L)|r_{K}(L)$ and $\rho_{K}(M,L)=\rho_{K}(M,L^{\prime})\rho_{L^{\prime}}(M,L)$ are not true in general. Let $M=\mathbb{Q}(\sqrt[8]{2}),L=\mathbb{Q}(\sqrt[4]{2}),L^{\prime}=\mathbb{Q}(\sqrt{2}),K=\mathbb{Q}$. Then $\rho_{K}(M,L)=r_{K}(L)=2\neq 4=2\cdot 2=r_{K}(L^{\prime})r_{L^{\prime}}(L)=\rho_{K}(M,L^{\prime})\rho_{L^{\prime}}(M,L)$.
We can generalize Cluster Magnification Theorem 3.3 for root capacity by using results from [2].
Theorem 3.6.
(Root Capacity Magnification Theorem) Consider $M/L/K$. Suppose $M^{\prime}/K$ and $L^{\prime}/K$ are obtained by strong cluster magnification from $M/K$ and $L/K$ respectively through the same $F/K$ with magnification factor $d$. Then $\rho_{K}(M^{\prime},L^{\prime})=d\ \rho_{K}(M,L)$.
Proof.
Let $\rho_{K}(M,L)=a\ r_{K}(L)$ and $\rho_{K}(M^{\prime},L^{\prime})=a^{\prime}\ r_{K}(L^{\prime})$ and $\rho_{F}(M^{\prime},L^{\prime})=a^{\prime\prime}\ r_{F}(L^{\prime})$ where $a,a^{\prime},a^{\prime\prime}$ are as in Definition 3.4. From Theorem 3.3, we have $r_{K}(L^{\prime})=d\ r_{K}(L)$.
By Definition 3.1, $M^{\prime}=MF$ and $L^{\prime}=LF$ where both the pairs $\tilde{M}$ and $F$ and $\tilde{L}$ and $F$ are linearly disjoint over $K$. By Lemma 8.1.7 in [2], $r_{F}(LF)=r_{K}(L)$. By Base Change Theorem for root capacity, Theorem 8.2.2 in [2], $\rho_{F}(MF,LF)=\rho_{K}(M,L)$. Hence $a=a^{\prime\prime}$.
Now by Corollary 8.1.5 in [2], $P/K$ is isomorphic to $LF/K$ $\iff$ $P/F$ is isomorphic to $LF/F$. Since $a^{\prime}$ is number of distinct fields inside $MF$ isomorphic to $LF$ over $K$ and $a^{\prime\prime}$ is number of distinct fields inside $MF$ isomorphic to $LF$ over $F$. Hence $a^{\prime}=a^{\prime\prime}$. Therefore $a=a^{\prime}$ and we are done. ∎
In [11], the notion of cluster tower of polynomials is introduced. See Section 5.2 in [2] for the group theoretic formulation. By using Section 3.2 in [2], we give the following field theoretic formulation.
Cluster tower of an extension: Consider $L/K$. Consider an ordering $(L_{1},L_{2},\dots,L_{s})$ of distinct fields isomorphic to $L$ over $K$ where $s=s_{K}(L)$. Now consider the following cluster tower of fields terminating at the Galois closure $\tilde{L}$.
$K\subseteq L_{1}\subseteq L_{1}L_{2}\subseteq\dots\subseteq L_{1}L_{2}\cdots L_{s}=\tilde{L}$.
The length of tower is number of distinct fields in the tower and the degrees of these distinct fields over $K$ form the degree sequence. Clearly length of tower $\leq s+1$.
Example 5.1.3 in [2] demonstrates that both the degree sequence and length of tower are dependent on the ordering of the $L_{i}$’s.
Proposition 3.7.
Suppose there exists a permutation $(i_{1},i_{2},\dots,i_{s})$ of $(1,2,\dots,s)$ such that
$K\subseteq L_{i_{1}}\subseteq L_{i_{1}}L_{i_{2}}\subseteq\dots\subseteq L_{i_{1}}L_{i_{2}}\cdots L_{i_{s}}=\tilde{L}$.
is a cluster tower for $L/K$ of length $s+1$. Then for each $0\leq a\leq s$ there exists an $M/K$ such that $\rho_{K}(M,L)=a\ r_{K}(L)$.
Proof.
For $a=0$, $M=K$ works. We claim that for $a\geq 1$, $M=L_{i_{1}}L_{i_{2}}\cdots L_{i_{a}}$ works. Since length of tower is $s+1$, we have each field in the tower to be a proper subset of successive field. Hence $\rho_{K}(L_{i_{1}}L_{i_{2}}\cdots L_{i_{a}},L)\geq a\ r_{K}(L)$. If $\rho_{K}(L_{i_{1}}L_{i_{2}}\cdots L_{i_{a}},L)>a\ r_{K}(L)$, then because of proper containment at each step, we will have $\rho_{K}(L_{i_{1}}L_{i_{2}}\cdots L_{i_{s}},L)>s\ r_{K}(L)=[L:K]$ which is a contradiction. Thus $\rho_{K}(M,L)=a\ r_{K}(L)$.∎
Proposition 3.8.
Let $M/K$ be obtained by strong cluster magnification from $L/K$ through $F/K$. Then
$K\subseteq L_{1}\subseteq L_{1}L_{2}\subseteq\dots\subseteq L_{1}L_{2}\cdots L_{s}$
is a cluster tower for $L/K$ of length $l$ if and only if
$K\subseteq L_{1}F\subseteq L_{1}L_{2}F\subseteq\dots\subseteq L_{1}L_{2}\cdots L_{s}F$
is a cluster tower for $M/K$ of length $l$. If degree sequence of first tower is $(a_{0},a_{1},\dots,a_{l-2})$, then degree sequence of second tower is $(a_{0}d,a_{1}d,\dots,a_{l-2}d)$ where $d=[F:K]$.
Proof.
By Corollary 8.1.5 in [2], the distinct fields inside $\bar{K}$ isomorphic to $M$ over $K$ are precisely $L_{i}F$ for $1\leq i\leq s$. Thus it is enough to show that, for any $1\leq k\leq s-1$, $L_{1}L_{2}\cdots L_{k}=L_{1}L_{2}\cdots L_{k}L_{k+1}$ if and only if $L_{1}L_{2}\cdots L_{k}F=L_{1}L_{2}\cdots L_{k}L_{k+1}F$.
Since $\tilde{L}\cap F=K$. Hence for $1\leq k\leq s$, we have $L_{1}L_{2}\cdots L_{k}\cap F=K$ and $L_{1}L_{2}\cdots L_{k}F\cap\tilde{L}=L_{1}L_{2}\cdots L_{k}$. If for any $1\leq k\leq s-1$, $L_{1}L_{2}\cdots L_{k}F=L_{1}L_{2}\cdots L_{k}L_{k+1}F$. Then $L_{1}L_{2}\cdots L_{k}=L_{1}L_{2}\cdots L_{k}F\cap\tilde{L}=L_{1}L_{2}\cdots L_{k}L_{k+1}F\cap\tilde{L}=L_{1}L_{2}\cdots L_{k}L_{k+1}$. We also have $L_{1}L_{2}\cdots L_{k}F:K]=[L_{1}L_{2}\cdots L_{k}:K][F:K]$.∎
4. Intersection Indicium : A Generalization of Ascending Index
The following is Theorem 9.0.4 in [2] referred to as the Ascending Index Magnification Theorem.
Theorem 4.1.
Let $M/K$ be obtained by strong cluster magnification from $L/K$ with magnification factor $d$. Then [M:K]=d[L:K][M:K]=d\ [L:K] and $t_{K}(M)=d\ t_{K}(L)$.
We define the following notion.
Definition 4.2.
Consider extensions $L/K$ and $M/K$. Let $L_{1},L_{2},\dots,L_{a}$ be all the distinct fields inside $M$ isomorphic to $L$ over $K$. Let $F^{\prime}=\cap_{i=1}^{a}L_{i}$. We define intersection indicium of $M$ with respect to $L$ (with base field $K$ fixed) as $\tau_{K}(M,L)=[F^{\prime}:K]$. If none of the fields isomorphic to $L/K$ is contained in $M$, then we define $\tau_{K}(M,L)=0$.
Remark 4.2.1.
Equivalently $\tau_{K}(M,L)=[G:H^{\prime}]$ where $G=Gal(\tilde{L}/K)$ (where $\tilde{L}$ is Galois closure of $L/K$) and $H^{\prime}$ is subgroup of $G$ generated by subgroups $H_{1},H_{2},\dots,H_{a}$ which fix $L_{1},L_{2},\dots,L_{a}$ respectively.
Remark 4.2.2.
Suppose $M/K$ and $M^{\prime}/K$ are isomorphic over $K$ and $L/K$ and $L^{\prime}/K$ are isomorphic over $K$. Then $\tau_{K}(M,L)=\tau_{K}(M^{\prime},L^{\prime})$ (If the isomorphism from $M$ to $M^{\prime}$ maps $L_{1},L_{2},\dots,L_{a}$ to $L_{i_{1}},L_{i_{2}},\dots,L_{i_{a}}$ respectively, then it maps $\cap_{j=1}^{a}L_{j}$ to $\cap_{j=1}^{a}L_{i_{j}}$).
The following are certain basic properties which demonstrate that intersection indicium is indeed a generalization of ascending index.
Proposition 4.3.
Consider an extension $L/K$.
  1. (1)
    $\tau_{K}(L,L)=[L:K]$. Also $\tau_{K}(L,K)=1$. If $L\neq K$ then $\tau_{K}(K,L)=0$.
  2. (2)
    $\tau_{K}(\tilde{L},L)=t_{K}(L)$ where $\tilde{L}$ is Galois closure of $L/K$.
  3. (3)
    For any extension $M/K$, $t_{K}(L)\ |\ \tau_{K}(M,L)$.
  4. (4)
    For any $M/K$, if $\tau_{K}(M,L)\neq 0$ then $\tau_{K}(M,L)\ |\ [L:K]$. Thus if $\tau_{K}(M,L)\neq 0$ then we have $\tau_{K}(M,L)=b\ t_{K}(L)$ where $b\ |\ u_{K}(L)$.
Proof.
  1. (1)
    Now $M=L$ contains only one field isomorphic to $L/K$ that is $L$. Thus $F^{\prime}=L$. Now if $L\neq K$ then $M=K$ does not contain any field isomorphic to $L/K$.
  2. (2)
    We have that $M=\tilde{L}$ contains all the $s_{K}(L)$ many distinct fields isomorphic to $L/K$. Thus $F^{\prime}=F$ by Proposition 2.2 (1) where $F/K$ is the unique intermediate extension of $L/K$ which is Galois with maximum possible degree. Since $t_{K}(L)=[F:K]$, we are done.
  3. (3)
    By Proposition 2.2 (1) we have that $F$ is the intersection of all the $s_{K}(L)$ many distinct fields isomorphic to $L/K$. Thus $F\subset F^{\prime}$. Hence $F:K]\ |\ [F^{\prime}:K]$.
  4. (4)
    Since $\tau_{K}(M,L)=[\cap_{i=1}^{a}L_{i}:K]$. Thus $\tau_{K}(M,L)$ divides $L_{1}:K]=[L:K]$. Since $L:K]=t_{K}(L)u_{K}(L)$ we are done.
The following connects the concepts of root capacity and intersection indicium.
Proposition 4.4.
Consider extensions $L/K$ and $M/K$.
  1. (1)
    $\rho_{K}(M,L)=r_{K}(L)\iff\tau_{K}(M,L)=[L:K]$.
  2. (2)
    $\rho_{K}(M,L)=[L:K]\implies\tau_{K}(M,L)=t_{K}(L)$. The converse is false as demonstrated in Example 4.5.
Example 4.5.
Consider $L/K$ which has Galois closure $\tilde{L}$ with Galois group as $\mathfrak{S}_{n}$ for $n>2$. By the argument in Example 2.3 (2), we have that intersection of any two distinct fields isomorphic to $L/K$ is $K$. Thus for an extension $M/K$, if $\tau_{K}(M,L)\neq 0$ or $[L:K]$ then $\tau_{K}(M,L)=1=t_{K}(L)$.
We also have that if $M$ is compositum of any two distinct fields isomorphic to $L/K$ then $\rho_{K}(M,L)=2=2\ r_{K}(L)$ by Example 3.5. Thus $\rho_{K}(M,L)\neq n=[L:K]$.
Some other properties which are easy to verify.
Proposition 4.6.
Consider an extension $L/K$.
  1. (1)
    For $M/M^{\prime}/K$ we have $\tau_{K}(M,L)\ |\ \tau_{K}(M^{\prime},L)$ if $\tau_{K}(M,L)\neq 0$.
  2. (2)
    For $K^{\prime}/K$ we have $\tau_{K}(M,L)\ |\ \tau_{K^{\prime}}(M,L)$.
We can generalize Ascending Index Magnification Theorem 4.1 for intersection indicium.
Theorem 4.7.
(Intersection Indicium Magnification Theorem) Consider $M/L/K$. Suppose $M^{\prime}/K$ and $L^{\prime}/K$ are obtained by strong cluster magnification from $M/K$ and $L/K$ respectively through the same $F/K$ with magnification factor $d$. Then $\tau_{K}(M^{\prime},L^{\prime})=d\ \tau_{K}(M,L)$.
Proof.
By proof of Theorem 3.6, the distinct fields inside $M^{\prime}$ isomorphic to $L^{\prime}/K$ are precisely $\{L_{i}F\}_{i=1}^{a}$ where $\{L_{i}\}_{i=1}^{a}$ are fields inside $M$ isomorphic to $L/K$. Let $G^{\prime}=Gal(\tilde{L^{\prime}}/K)$ and $G=Gal(\tilde{L}/K)$ and $R=Gal(F/K)$. By Proposition 4.1.5 (3) in [2], we can identify $G^{\prime}$ with $G\times R$. So $Gal(\tilde{L^{\prime}}/\tilde{L})=1\times R$. If $H_{i}=Gal(\tilde{L}/L_{i})$ then $H_{i}\times R=Gal(\tilde{L^{\prime}}/L_{i})$. Also $G\times 1=Gal(\tilde{L^{\prime}}/F)$. Thus $(H_{i}\times R)\cap(G\times 1)=H_{i}\times 1=Gal(\tilde{L^{\prime}}/(L_{i}F))$. Hence $Gal(\tilde{L^{\prime}}/(\cap_{i=1}^{a}L_{i}F))$ is generated by subgroups $\{H_{i}\times 1\}_{i=1}^{a}$ which is $H^{\prime}\times 1$ where $H^{\prime}$ is subgroup of $G$ generated by subgroups $\{H_{i}\}_{i=1}^{a}$. Also $Gal(\tilde{L^{\prime}}/(\cap_{i=1}^{a}L_{i}))=H^{\prime}\times R$. Since $(H^{\prime}\times R)\cap(G\times 1)=H^{\prime}\times 1$. Hence $(\cap_{i=1}^{a}L_{i})F=(\cap_{i=1}^{a}L_{i}F)$. Since $\tilde{L}\cap F=K$ we have $(\cap_{i=1}^{a}L_{i})\cap F=K$ where $F/K$ is Galois. Thus we have $\tau_{K}(M^{\prime},L^{\prime})=[(\cap_{i=1}^{a}L_{i}F):K]=[(\cap_{i=1}^{a}L_{i})F:K]=[(\cap_{i=1}^{a}L_{i}):K][F:K]=\tau_{K}(M,L)\ d$. ∎
Proposition 4.8.
Consider the case in Example 2.3 (1) for $n$ odd. Let $L=\mathbb{Q}(a)$ and $M=\mathbb{Q}(a,\zeta^{l})$ for an $l|n$. Then $\rho_{\mathbb{Q}}(M,L)=n/l$ and $\tau_{\mathbb{Q}}(M,L)=l$ and thus $\tau_{\mathbb{Q}}(M,L)\rho_{\mathbb{Q}}(M,L)=[L:\mathbb{Q}]$.
Proof.
Suppose $\mathbb{Q}(a\zeta^{i})\subset M$. Then $\zeta^{i}\in\mathbb{Q}(a,\zeta^{l})$ and so $\zeta^{i}\in\mathbb{Q}(a,\zeta^{l})\cap\mathbb{Q}(\zeta)$. Since $\mathbb{Q}(a)\cap\mathbb{Q}(\zeta)=\mathbb{Q}$. Thus $\mathbb{Q}(a,\zeta^{l})\cap\mathbb{Q}(\zeta)=\mathbb{Q}(\zeta^{l})$ and so $\zeta^{i}\in\mathbb{Q}(\zeta^{l})$. Let $d=gcd(i,l)$. Then $\zeta^{d}\in\mathbb{Q}(\zeta^{l})$. Now we have $\mathbb{Q}(\zeta^{l}):\mathbb{Q}]=\phi(n/l)$ where $\phi$ is the Euler totient function. Since $d|l$, we have $d|n$. Thus $\mathbb{Q}(\zeta^{d}):\mathbb{Q}]=\phi(n/d)$. Since $\mathbb{Q}(\zeta^{d})\subset\mathbb{Q}(\zeta^{l})$. Thus $\phi(n/d)\ |\ \phi(n/l)$. Since $d|l$ we have $(n/l)|(n/d)$. Thus $\phi(n/l)\ |\ \phi(n/d)$ which implies that $\phi(n/d)=\phi(n/l)$. Since $\phi(n/d)=\phi((n/d)/(l/d))$ and $n$ is odd. Thus $n/d=n/l$ that is $d=l$. Thus we have $l|i$ (Refer Proposition 5.1.2 in [2] for a list of properties of $\phi$ that we have used here).
Thus the distinct fields inside $M$ isomorphic to $L/\mathbb{Q}$ are precisely $\{\mathbb{Q}(a\zeta^{yl})\}_{y=0}^{(n/l-1)}$. Since $n$ is odd, we have $r_{\mathbb{Q}}(L)=1$. Thus $\rho_{\mathbb{Q}}(M,L)=n/l$.
Now let $F^{\prime}$ be intersection of all these distinct fields inside $M$ isomorphic to $L/\mathbb{Q}$. Then we have $\tau_{\mathbb{Q}}(L_{i},L)=[F^{\prime}:K]$. Clearly $a^{n/l}\in F^{\prime}$. We will show that $F^{\prime}=\mathbb{Q}(a^{n/l})$. Since $l|n$, we will have $F^{\prime}:K]=l$.
Suppose $\gamma\in F^{\prime}$. So $\gamma\in\mathbb{Q}(a)\cap\mathbb{Q}(a\zeta^{l})$. Hence $\gamma=a_{0}+a_{1}a+\dots+a_{n-1}a^{n-1}=b_{0}+b_{1}(a\zeta^{l})+\dots+b_{n-1}(a\zeta^{l})^{n-1}$ for $a_{j},b_{j}\in\mathbb{Q}$ for all $0\leq j\leq n-1$. By arguments as in Example 2.3 (1), we have for all $0\leq j\leq n-1$ that $a_{j}=b_{j}(\zeta^{l})^{j}$. Since $n$ is odd, $a_{j}\neq 0$ implies n|(lj)n\ |\ (lj) that is (n/l)|j(n/l)\ |\ j. Thus if $a_{j}\neq 0$ then $a^{j}\in\mathbb{Q}(a^{n/l})$ and thus $\gamma\in\mathbb{Q}(a^{n/l})$.∎
Remark 4.8.1.
Proposition 4.8 is a generalization of Theorem 3.4 (3) in [10].
Remark 4.8.2.
For extensions $L/K$ and $M/K$, the statement $\tau_{K}(M,L)\rho_{K}(M,L)=[L:K]$ is not true in general which is demonstrated by the case in Example 4.5.
5. Inverse Root Capacity Problem
The following is Theorem 3.1.1 in [2] referred to as the Inverse cluster size problem for number fields.
Theorem 5.1.
Let $K$ be a number field. Let $n>2$ and $r|n$. Then there exists an extension $L/K$ of degree $n$ with cluster size $r$ with Galois closure having solvable Galois group.
We establish the following in this section as a generalization of Theorem 5.1.
Theorem 5.2.
(Inverse Root Capacity Problem for Number Fields) Let $K$ be a number field. Given $(n,r,\rho)$ where $n>2$ and $r|n$ and $r|\rho$ and $\rho\neq n-1$. There exist extensions $L/K$ and $M/K$ such that $[L:K]=n$ and $r_{K}(L)=r$ and $\rho_{K}(M,L)=\rho$. For $\rho\neq 0$, we get $M/K$ as an extension of $L/K$ contained in $\tilde{L}$.
We discuss two approaches to prove the above result. The first one uses Theorem 3.6 but excludes certain cases. The second approach gives a complete answer.
Remark 5.2.1.
Assuming the condition $\rho\neq n-1$ is necessary in Theorem 5.2. Suppose $L/K$ is degree $n$ extension. Then there doesn’t exist $M/K$ such that $\rho_{K}(M,L)=n-1$. Assume on the contrary that such $M/K$ exists. Since $r_{K}(L)|\rho_{K}(M,L)$ and $r_{K}(L)|n$. Thus $r_{K}(L)=1$ and $s_{K}(L)=n$. Let $L=K(\alpha)$ and $f$ be minimal polynomial for $\alpha$ over $K$. Since $\rho_{K}(M,L)=n-1$, $M$ contains n1n-1 roots of $f$. Since sum of all roots of $f$ lies in $K$. Hence $M$ contains the $n$-th root as well which is a contradiction.
By results in [18] on hilbertian fields and by the final proposition in [14], we have the following.
Lemma 5.3.
Let $K$ be a number field and $n\in\mathbb{N}$. Then we have ${\mathfrak{S}}_{n}$ to be realizable as a Galois group for infinitely many pairwise linearly disjoint Galois extensions over $K$. Thus there exist infinitely many pairwise linearly disjoint extensions over $K$ of degree $n$ with Galois closures having Galois group ${\mathfrak{S}}_{n}$ over $K$.
Lemma 5.4 (Lemma 3, [11]).
Let $K$ be a number field and $L/K$ be a finite extension. For any positive integer $d\geq 2$, there exist infinitely many cyclic Galois extensions $L^{\prime}/K$ of degree $d$ such that $L$ and $L^{\prime}$ are linearly disjoint over $K$.
The First Approach: This excludes the cases (1) $n=2r$ and (2) $\rho=n-r$ for $n>2r$.
Proof.
Let $n/r=s$ and $\rho/r=a$ and let $s\neq 2$ and $a\neq s-1$. By Lemma 5.3, there exists an irreducible polynomial $f$ over $K$ of degree $s$ with Galois group ${\mathfrak{S}}_{s}$. This $f$ satisfies $r_{K}(f)=1$. Let roots of $f$ be $\alpha_{i}$ for $1\leq i\leq s$. For $1\leq k\leq s-1$, let $L_{k}=K(\alpha_{1},\dots,\alpha_{k})$. As noted in Example 6.2.7 in [2], $r_{K}(L_{1})=1$ and $L_{1}:K]=s$ and $L_{s-1}$ is Galois closure of $L_{1}/K$ and $\rho_{K}(L_{k},L_{1})=k$ for $1\leq k\leq(s-2)$ and $\rho_{K}(L_{s-1},L_{1})=s$.
By Lemma 5.4, there exists $F/K$ Galois of degree $r$ such that $L_{s-1}$ and $F$ are linearly disjoint over $K$. Let $L=L_{1}F$. By Theorem 3.3, $[L:K]=rs=n$ and $r_{K}(L)=r$. Then for $a=0$, $M=K$ works. Let $M=L_{a}F$ for $a\neq 0,s$ and let $M=L_{s-1}F$ for $a=s$. By Theorem 3.6, $\rho_{K}(M,L)=r\ \rho_{K}(L_{a},L_{1})=ar=\rho$ for $a\neq 0,s$ and $\rho_{K}(M,L)=r\ \rho_{K}(L_{s-1},L_{1})=sr=\rho$ for $a=s$. ∎
The Second Approach: This gives a complete answer.
Proof.
For $r=1$ we use the first approach itself. For $r>1$, we proceed as we did in the proof of Theorem 5.1 in [2]. Thus we have that there exists an extension $L/K$ with Galois closure $\tilde{L}$ such that $G=Gal(\tilde{L}/K)$ is solvable such that action of the group is transitive on $n$ points, and a point stabiliser fixes precisely $r$ points. We have $G=({\mathbb{Z}}/r{\mathbb{Z}})^{s}\rtimes{\mathbb{Z}}/s{\mathbb{Z}}$ with semidirect product group law
$((a_{1},\dots,a_{s}),b)\cdot((c_{1},\dots,c_{s}),d)=((a_{1},\dots,a_{s})+(b\cdot(c_{1},\dots,c_{s})),b+d),$
where $b\cdot(c_{1},\dots,c_{s})=(c_{b+1},\dots,c_{s},c_{1},\dots,c_{b})$ for $b\neq 0$ & $0\cdot(c_{1},\dots,c_{s})=(c_{1},\dots,c_{s})$.
One can verify that $((a_{1},\dots,a_{s}),b)^{-1}=((-a_{s-b+1},\dots,-a_{s},-a_{1},\dots,-a_{s-b}),-b)$ and
$((a_{1},\dots,a_{s}),b)\cdot((c_{1},\dots,c_{s-1},0),0)\cdot((a_{1},\dots,a_{s}),b)^{-1}=((c_{b+1},\dots,c_{s-1},0,c_{1},\dots,c_{b}),0).$
Any point stabiliser is isomorphic to $({\mathbb{Z}}/r{\mathbb{Z}})^{s-1}$. We have $[L:K]=n$ and $r_{K}(L)=r$. The $s=s_{K}(L)$ many subgroups of $G$ fixing the $s$ many distinct fields $L_{i}$’s isomorphic to $L/K$ are $H_{i}=(({\mathbb{Z}}/r{\mathbb{Z}})^{i-1}\times 0\times({\mathbb{Z}}/r{\mathbb{Z}})^{s-i})\times 0$ for $1\leq i\leq s$. Observe that
$G\supsetneq H_{1}\supsetneq H_{1}\cap H_{2}\supsetneq\dots\supsetneq H_{1}\cap H_{2}\cap\cdots\cap H_{s}=0.$
Thus by the group theoretic formulation for cluster towers in Section 5.2 in [2], we have
$K\subseteq L_{1}\subseteq L_{1}L_{2}\subseteq\dots\subseteq L_{1}L_{2}\cdots L_{s}=\tilde{L}$.
is a cluster tower for $L/K$ of length $s+1$. Hence we are done by Proposition 3.7. ∎
Remark 5.4.1.
Consider the case in the first approach above. Now $L/K$ is obtained by strong cluster magnification from $K(\alpha_{1})/K$ through $F/K$. Thus from the alternate proof of cluster magnification theorem in Section 8 in [2], we have $s_{K}(L)=s_{K}(K(\alpha_{1}))=s$ and $K(\alpha_{i})F$ for $1\leq i\leq s$ are the $s$ many distinct fields isomorphic to $L$ over $K$. Hence we have by Proposition 3.8 that
$K\subseteq L_{1}F\subseteq L_{2}F\subseteq\dots\subseteq L_{s-1}F=\tilde{L}$.
is a cluster tower for $L/K$ of length $s$ where $L_{k}=K(\alpha_{1},\dots,\alpha_{k})$ for $1\leq k\leq s-1$. The degree sequence of the cluster tower is $(n,n\ ^{(s-1)}P_{1},n\ ^{(s-1)}P_{2},\dots,n\ ^{(s-1)}P_{(s-2)})$. The degree sequence and length of tower are independent of the ordering of the $K(\alpha_{i})F$’s.
Remark 5.4.2.
In the second approach above we can compute the degree sequence of the considered cluster tower for $L/K$ which turns out to be $(n,nr,nr^{2},\dots,nr^{s-1})$. It is easy to see that degree sequence and length of tower are independent of the ordering of the $L_{i}$’s.
Remark 5.4.3.
For $1<r\leq n/3$ and $r|n$, consider the field $L/K$ in second approach above. Thus $3\leq s=n/r<n$. For $1\leq a\leq s-1$, let $M_{a}=L_{1}L_{2}\cdots L_{a}$. Now Galois closure of $M_{a}/K$ is $\tilde{L}$. Thus for $2\leq a\leq s-1$, we have that $M_{a}/K$ is not obtained by strong cluster magnification from $M_{1}/K$.
Now for $1\leq a\leq s-1$, the subgroup of $G$ fixing $M_{a}$ is $G_{a}=H_{1}\cap H_{2}\cap\cdots\cap H_{a}$. We have the notions of unique descending chains and unique ascending chains for extensions introduced in Section 7 in [2]. By using the group law we can show that for $1\leq a\leq s-1$, $N_{G}(G_{a})=G_{a}^{G}=({\mathbb{Z}}/r{\mathbb{Z}})^{s}\times\{0\}$ where $G_{a}^{G}$ is the normal closure of $G_{a}$ in $G$.
Hence for any $1\leq a\leq s-1$ the unique descending chain for $M_{a}/K$ is $M_{a}\supsetneq\tilde{L}^{({\mathbb{Z}}/r{\mathbb{Z}})^{s}\times\{0\}}\supsetneq K$ and the unique ascending chain for $M_{a}/K$ is $K\subsetneq\tilde{L}^{({\mathbb{Z}}/r{\mathbb{Z}})^{s}\times\{0\}}\subsetneq M_{a}$ and both the unique chains coincide. Thus $r_{K}(M_{a})=[N_{G}(G_{a}):G_{a}]=[G_{a}^{G}:G_{a}]=u_{K}(M_{a})=r^{s}/r^{s-a}=r^{a}$ and
$s_{K}(M_{a})=[G:N_{G}(G_{a})]=[G:G_{a}^{G}]=t_{K}(M_{a})=s$.
Thus we have for $2\leq a\leq s-1$ that $M_{a}/K$ and the subextension $M_{1}/K$ serve as a counterexample for the converse of Theorem 8.4.1 in [2].
Remark 5.4.4.
For extensions $L/K$ and $M/K$ with $\rho_{K}(M,L)\neq 0$, it is not necessary that we have $\rho_{K}(M,L)\ |\ [M:K]$. Thus it is not even necessary that $\rho_{K}(M,L)\ |\ [L:K]$.
For example consider the case in Remark 5.4.3. Then $\rho_{K}(M_{a},L)=a\ r_{K}(L)=ar$ and $M_{a}:K]=r^{a}\ s$. Thus $\rho_{K}(M_{a},L)\ |\ [M_{a}:K]$ if and only if $a|r^{a-1}s$ which is not always true.
6. Inverse Intersection Indicium Problem
The following is Theorem 9.0.5 in [2] which is referred to as the Inverse ascending index problem for number fields.
Theorem 6.1.
Let $K$ be a number field. Let $n>2$ and $t|n$. Then there exists an extension $L/K$ of degree $n$ with ascending index $t_{K}(L)=t$ with Galois closure having solvable Galois group.
We establish the following in this section as a generalization of Theorem 6.1.
Theorem 6.2.
(Inverse Intersection Indicium Problem for Number Fields) Let $K$ be a number field. Given $(n,t,\tau)$ where $n>2$ and $t\ |\ \tau\ |\ n$ (except for the cases listed in Remark 6.6.1). There exist extensions $L/K$ and $M/K$ such that $[L:K]=n$ and $t_{K}(L)=t$ and $\tau_{K}(M,L)=\tau$. We get $M/K$ as an extension of $L/K$ contained in $\tilde{L}$.
Example 6.3 below demonstrates that we can’t directly use the second approach used for proving Inverse root capacity problem Theorem 5.2 to solve the above problem.
Example 6.3.
Let $K$ be a number field and $n>2$ and $t|n$ with $t>1$. Let $r=n/t$ for $t<n$ and $r=n$ for $t=n$. By the proof of Theorem 6.1 in [2], we have $L/K$ with $[L:K]=n$ and $t_{K}(L)=t$ and $r_{K}(L)=r$ which is the same as the field extension $L/K$ in second approach of proof of Theorem 5.2. Consider the notations as in that proof above.
If $L_{i}$ and $L_{j}$ are two distinct fields isomorphic to $L/K$ then the subgroups $H_{i}$ and $H_{j}$ fixing them clearly generate the subgroup $({\mathbb{Z}}/r{\mathbb{Z}})^{s}\times\{0\}$ which is same as $H^{G}$ as shown in Theorem 7.3.4 in [2]. Thus for an extension $M/K$, if $\tau_{K}(M,L)\neq 0$ or $[L:K]$ then $\tau_{K}(M,L)=t=t_{K}(L)$. In particular, for $L/K$ non-Galois ($t<n$) and $2\leq a\leq s$, we have $\tau_{K}(L_{1}L_{2}\cdots L_{a},L)=t=t_{K}(L)$.
Similarly one can see in the light of Example 4.5, that the first approach of Theorem 5.2 is also not directly useful to solve the problem.
We would need the following result which is a slightly generalized version of Theorem 4.7. The hypothesis is inspired from Lemma 2 and the proof of the Main Theorem in [12].
Lemma 6.4.
Consider $M/L/K$. Let $J/K$ be an extension such that $\tilde{M}\cap\tilde{J}=K$. Then we have $[LJ:K]=[L:K][J:K]$ and $t_{K}(LJ)=t_{K}(L)t_{K}(J)$ and $\rho_{K}(MJ,LJ)=r_{K}(J)\rho_{K}(M,L)$ and $\tau_{K}(MJ,LJ)=[J:K]\ \tau_{K}(M,L)$.
Proof.
Since $\tilde{M}\cap\tilde{J}=K$, we have $\tilde{L}\cap\tilde{J}=K$. Let $G={{\rm Gal}}(\tilde{L}/K)$, $H={{\rm Gal}}(\tilde{L}/L)$, $S={{\rm Gal}}(\tilde{J}/K)$ and $T={{\rm Gal}}(\tilde{J}/J)$. Let $G^{\prime}=Gal(\tilde{L}\tilde{J}/K)$ and $H^{\prime}=Gal(\tilde{L}\tilde{J}/LJ)$. We can identify $G^{\prime}$ with $G\times S$. Thus $H^{\prime}=(H\times S)\cap(G\times T)=H\times T$ and hence $[LJ:K]=[L:K][J:K]$.
Also $N_{G^{\prime}}(H^{\prime})=N_{G}(H)\times N_{S}(T)$. Thus $s_{K}(LJ)=[G^{\prime}:N_{G^{\prime}}(H^{\prime})]=s_{K}(L)s_{K}(J)$. Thus the $s_{K}(LJ)$ many distinct fields isomorphic to $LJ/K$ are precisely $L_{i}J_{j}$ for $1\leq i\leq s_{K}(L)$ and $1\leq j\leq s_{K}(J)$ where $L_{i}$’s are all the distinct fields isomorphic to $L/K$ and $J_{j}$’s are all the distinct fields isomorphic to $J/K$. We also have $r_{K}(LJ)=r_{K}(L)r_{K}(J)$. We can also show that $H^{\prime G^{\prime}}=H^{G}\times T^{S}$. Hence $t_{K}(LJ)=[G^{\prime}:H^{\prime G^{\prime}}]=t_{K}(L)t_{K}(J)$.
Since $\tilde{M}\cap\tilde{J}=K$. Hence $MJ\cap\tilde{J}=J$ and $MJ\cap\tilde{M}=M$. Now suppose $L_{i}J_{j}\subset MJ$. Thus $J_{j}\subset MJ$. Hence $J_{j}\cap\tilde{J}\subset MJ\cap\tilde{J}$. Therefore $J_{j}\subset J$. Hence $J_{j}=J$. Now $LiJ\subset MJ$. Thus $L_{i}J\cap\tilde{M}\subset MJ\cap\tilde{M}=M$. Since $L_{i}\subset\tilde{M}$ and $\tilde{M}\cap\tilde{J}=K$, we have $L_{i}J\cap\tilde{M}=L_{i}$. Hence $L_{i}\subset M$. Thus the distinct fields inside $MJ$ isomorphic to $LJ/K$ are precisely $\{L_{i}J\}_{i=1}^{a}$ where $\{L_{i}\}_{i=1}^{a}$ are fields inside $M$ isomorphic to $L/K$. Hence $\rho_{K}(MJ,LJ)/r_{K}(LJ)=\rho_{K}(M,L)/r_{K}(L)$. Thus $\rho_{K}(MJ,LJ)=r_{K}(J)\rho_{K}(M,L)$.
[…]
Lemma 6.5 (Lemma 3.1.4, [2]).
Let $K$ be a perfect field and $K^{\prime}/K$ be a finite extension. If for a group $G$, direct products $G^{n}$ are realisable as Galois group over $K$ for each $n\in\mathbb{N}$ then they are realisable as Galois groups over the field $K^{\prime}$.
Lemma 6.6.
Given $L/K$.
  1. (1)
    If for a group $G$, direct products $G^{n}$ are realisable as Galois group over $K$ for each $n\in\mathbb{N}$ then we have a Galois extension $L^{\prime}/K$ such that $Gal(L^{\prime}/K)=G$ and $L^{\prime}\cap L=K$ that is $L^{\prime}$ and $L$ are linearly disjoint over $K$.
  2. (2)
    If $K$ is a number field and $G$ is a solvable group then we have a Galois extension $L^{\prime}/K$ such that $Gal(L^{\prime}/K)=G$ and $L^{\prime}\cap L=K$.
  3. (3)
    If $K$ is a number field then for any $n\in\mathbb{N}$, we have a degree $n$ extension $L^{\prime}/K$ such that $Gal(\tilde{L^{\prime}}/K)=\mathfrak{S}_{n}$ and $L^{\prime}\cap L=K$.
Proof.
  1. (1)
    Follows from the proof in [2] of Lemma 6.5 (essentially the proof of Theorem 4.2 in [3]).
  2. (2)
    Since direct product of solvable groups is solvable, the $n$-fold direct products $G^{n}$ for $n\in\mathbb{N}$ are solvable. By Shafarevich’s theorem ([16]), $G^{n}$ for $n\in\mathbb{N}$ are realizable as Galois groups over $\mathbb{Q}$. By Lemma 6.5 they are realisable over number field $K$. By part (1) we are done.
  3. (3)
    The proof is similar to proof of Part (a) of Main theorem in [12] but we provide it nonetheless. By Lemma 5.3, there exist infinitely many degree $n$ extensions over $K$ with Galois closures having Galois group ${\mathfrak{S}}_{n}$.
    Since $L/K$ is separable, it has finitely many intermediate fields. Thus there is a degree $n$ extension $L^{\prime}/K$ such that $L^{\prime}\not\subset L$ and $Gal(\tilde{L^{\prime}}/K)=\mathfrak{S}_{n}$. Now $L^{\prime}\cap L$ is an intermediate field for $L^{\prime}/K$. Since $\mathfrak{S}_{n-1}$ is a maximal subgroup of $\mathfrak{S}_{n}$, we have $L^{\prime}\cap L=K$ or $L^{\prime}\cap L=L^{\prime}$. The latter case is not possible since $L^{\prime}\not\subset L$. Hence $L^{\prime}\cap L=K$.
Now we prove Theorem 6.2 excluding certain cases which are listed in Remark 6.6.1.
Proof.
Case 1 : Suppose $\tau=t$ or $\tau=n$. Thus by Theorem 6.1 for $n>2$ with $t|n$, we have an extension $L/K$ with $[L:K]=n$ and $t_{K}(L)=t$. Thus for $M=\tilde{L}$ we have $\tau_{K}(M,L)=t_{K}(L)=t$ and for $M=L$ we have $\tau_{K}(M,L)=[L:K]=n$.
Case 2 : Suppose $n/\tau>2$ and $\tau>2$. Then by Lemma 5.3, we have an extension $L^{\prime}/K$ of degree $n/\tau$ such that its Galois closure $\tilde{L^{\prime}}$ has Galois group $\mathfrak{S}_{n/\tau}$ over $K$. Thus by Theorem 7.3.6 (2) in [2] we have $t_{K}(L^{\prime})=1$ and $\tau_{K}(\tilde{L^{\prime}},L^{\prime})=t_{K}(L^{\prime})=1$. Since we have $t|\tau$ and $\tau>2$. Thus by Theorem 6.1, we have $J/K$ with $J:K]=\tau$ and $t_{K}(J)=t$. Now by Lemma 6.6 we can choose the fields such that $\tilde{L^{\prime}}\cap\tilde{J}=K$. Let $L=L^{\prime}J$ and $M=\tilde{L^{\prime}}J$. Thus by Lemma 6.4 we have $L:K]=[L^{\prime}:K][J:K]=(n/\tau)\cdot\tau=n$ and $t_{K}(L)=t_{K}(L^{\prime})t_{K}(J)=1\cdot t=t$ and $\tau_{K}(M,L)=[J:K]\tau_{K}(\tilde{L^{\prime}},L^{\prime})=\tau\cdot 1=\tau$.
Case 3 : Suppose $n=2\tau$ and $t>1$ and $\tau/t>2$. Then we have an extension $J/K$ of degree $\tau/t$ such that its Galois closure $\tilde{J}$ has Galois group $\mathfrak{S}_{\tau/t}$ over $K$. Thus $t_{K}(J)=1$. Since we have $t|2t$ and $t>1$. Thus by Theorem 6.1, we have $L^{\prime}/K$ with $L^{\prime}:K]=2t$ and $t_{K}(L^{\prime})=t$. Also $\tau_{K}(\tilde{L^{\prime}},L)=t$. Now by Lemma 6.6 we can choose the fields such that $\tilde{L^{\prime}}\cap\tilde{J}=K$. Let $L=L^{\prime}J$ and $M=\tilde{L^{\prime}}J$. Thus by Lemma 6.4 we have $L:K]=[L^{\prime}:K][J:K]=(2t)\cdot(\tau/t)=2\tau=n$ and $t_{K}(L)=t_{K}(L^{\prime})t_{K}(J)=t\cdot 1=t$ and $\tau_{K}(M,L)=[J:K]\tau_{K}(\tilde{L^{\prime}},L^{\prime})=(\tau/t)\cdot t=\tau$.
Case 4 : Suppose $n=2\tau$ and $\tau=2t$ and $t=2m$ with $m>1$. So $n=8m$ and $\tau=4m$. By Theorem 6.1, we have extension $L^{\prime}/K$ with $L^{\prime}:K]=4$ and $t_{K}(L^{\prime})=2$ and extension $J/K$ with $[J:K]=2m$ and $t_{K}(J)=m$. Now by Lemma 6.6 we can choose the fields such that $\tilde{L^{\prime}}\cap\tilde{J}=K$. Let $L=L^{\prime}J$ and $M=\tilde{L^{\prime}}J$. Thus by Lemma 6.4 we have $L:K]=[L^{\prime}:K][J:K]=4\cdot 2m=8m=n$ and $t_{K}(L)=t_{K}(L^{\prime})t_{K}(J)=2\cdot m=2m=t$ and $\tau_{K}(M,L)=[J:K]\tau_{K}(\tilde{L^{\prime}},L^{\prime})=2m\cdot 2=4m=\tau$.
Case 5 : $n=8,\tau=4,t=2$. First let $K=\mathbb{Q}$. Consider the polynomial $x^{8}-3$ which is irreducible over $\mathbb{Q}$ having a solvable Galois group by results in [8]. Let $a=3^{1/8}$ be the positive real root of the polynomial and $\iota$ be a primitive $4$-th root of unity. Let $L=\mathbb{Q}(a)$ and $M=\mathbb{Q}(a,\iota)$. Thus $\tau_{\mathbb{Q}}(M,L)=[\mathbb{Q}(a)\cap\mathbb{Q}(ai):\mathbb{Q}]=[\mathbb{Q}(a^{2}):\mathbb{Q}]=4$. Also by Example 2.3 (1), $t_{\mathbb{Q}}(L)=[\mathbb{Q}(a^{4}):\mathbb{Q}]=2$. Since the Galois group of Galois closure $\tilde{L}$ over $\mathbb{Q}$ is solvable and $M\subset\tilde{L}$, we are done for any number field $K$ by Lemma 6.6 (2) and Remark 4.2.1.
Remark 6.6.1.
The following cases for $(n,t,\tau)$ are not covered by the above proof.
  1. (1)
    $(n,1,2)$ for $n>2$ and $2|n$.
  2. (2)
    $(2\tau,1,\tau)$ for $\tau>1$.
  3. (3)
    $(4t,t,2t)$ for $t\geq 1$ odd.
Note that the subcase $(4,1,2)$ is common to all three classes in the above remark.
Proposition 6.7.
We don’t have a positive answer for Theorem 6.2 for the case $(n,t,\tau)=(4,1,2)$ for any number field.
Proof.
Let $K$ be any number field. Supppose we have extensions $L/K$ and $M/K$ such that
$[L:K]=4$, $t_{K}(L)=1$ and $\tau_{K}(M,L)=2$. The groups which can possibly occur as Galois group $G=Gal(\tilde{L}/K)$ are $\mathfrak{S}_{4},\mathfrak{A}_{4},D_{4},C_{4}$ and $C_{2}^{2}$. Now since $t_{K}(L)\neq 4$, $G$ cannot be abelian. Thus $G\neq C_{4},C_{2}^{2}$. It is easy to see that if $G=D_{4}$ then $t_{K}(L)=2$ which is not the case. Finally $G\neq\mathfrak{S}_{4},\mathfrak{A}_{4}$ otherwise $\tau_{K}(M,L)\neq 2$ since $\mathfrak{S}_{3}$ and $\mathfrak{A}_{3}$ are maximal subgroups in $\mathfrak{S}_{4}$ and $\mathfrak{A}_{4}$ respectively. ∎
7. Minimal Generating Sets of Galois Closure
We have the notion of minimal generating sets of the splitting field of a polynomial introduced by the author and Vanchinathan in Section 2 of their work in [10]. The following is a field theoretic formulation of the same in light of Proposition 2.2 in [10].
Consider an extension $L/K$. Let $S=\{L_{i}\}_{i=1}^{s}$ where $L_{i}$’s are the $s=s_{K}(L)$ many distinct fields isomorphic to $L$ over $K$. For any set $B\subset S$ we denote compositum of fields in $B$ as $L_{B}$.
Definition 7.1.
A set $B\subset S$ is called a minimal generating set of the Galois Closure $\tilde{L}$ of $L/K$ if the following hold.
  1. (1)
    $L_{B}=\tilde{L}$
  2. (2)
    For any set $A\subsetneq B$, we have $L_{A}\neq\tilde{L}$.
Remark 7.1.1.
Equivalently $B=\{L_{i_{j}}\}_{j=1}^{m}\subset S$ is a minimal generating set of the Galois closure of $L/K$ if $\cap_{j\in[m]}\ H_{i_{j}}=1$ and for any $J\subsetneq[m]$ we have $\cap_{j\in J}\ H_{i_{j}}\neq 1$ where $m]=\{1,2,\dots,m\}$ and $G=Gal(\tilde{L}/K)$ (where $\tilde{L}$ is Galois closure of $L/K$) and $H_{i_{j}}$’s are subgroups of $G$ which fix $L_{i_{j}}$’s respectively (which are conjugates of $Gal(\tilde{L}/L)$ in $G$).
The following is a reformulation of Proposition 2.4 in [10].
Lemma 7.2.
Let $C\subset S$ be a generating set of the Galois closure of $L/K$ i.e. $L_{C}=\tilde{L}$. Then there exists $B\subset C$ such that $B$ is a minimal generating set of the Galois closure of $\tilde{L}$.
The following is a reformulation of Theorem 3.1 (1) in [10].
Lemma 7.3.
Consider $B=\{L_{i_{j}}\}_{j=1}^{m}\subset S$. We have that $B$ is a minimal generating set of the Galois closure of $L/K$ if and only if for every permutation $(l_{1},l_{2},\dots,l_{m})$ of $(i_{1},i_{2},\dots,i_{m})$,
$K\subseteq L_{l_{1}}\subseteq L_{l_{1}}L_{l_{2}}\subseteq\dots\subseteq L_{l_{1}}L_{l_{2}}\cdots L_{l_{m}}$
is a cluster tower for $L/K$ of length $m+1$.
In this case we refer any such cluster tower as a minimal cluster tower.
Proposition 7.4.
Suppose $M/K$ is obtained by strong cluster magnification from $L/K$ through $F/K$. Then $B=\{L_{i_{j}}\}_{j=1}^{m}\subset S$ is a minimal generating set of the Galois closure of $L/K$ if and only if $B^{\prime}=\{L_{i_{j}}F\}_{j=1}^{m}$ is a minimal generating set of the Galois closure of $M/K$.
Proof.
Suppose $B$ is a minimal generating set of the Galois closure of $L/K$. Thus by Lemma 7.3, this is equivalent to saying that for every permutation $(l_{1},l_{2},\dots,l_{m})$ of $(i_{1},i_{2},\dots,i_{m})$,
$K\subseteq L_{l_{1}}\subseteq L_{l_{1}}L_{l_{2}}\subseteq\dots\subseteq L_{l_{1}}L_{l_{2}}\cdots L_{l_{m}}$
is a cluster tower for $L/K$ of length $m+1$. Hence by Proposition 3.8, this is equivalent to saying that for every permutation $(l_{1},l_{2},\dots,l_{m})$ of $(i_{1},i_{2},\dots,i_{m})$,
$K\subseteq L_{l_{1}}F\subseteq L_{l_{1}}L_{l_{2}}F\subseteq\dots\subseteq L_{l_{1}}L_{l_{2}}\cdots L_{l_{m}}F$
is a cluster tower for $M/K$ of length $m+1$. Therefore by Lemma 7.3, this is equivalent to $B^{\prime}$ being a minimal generating set of the Galois closure of M/KM/K.∎
Theorem 2.5 in [10] demonstrates that two minimal generating sets of Galois closure need not have same cardinalities. Thus we also have the following notion.
Definition 7.5.
$B\subset S$ is said to be minimum minimal generating set of the Galois closure of $L/K$ if $B$ is minimal generating set with least possible cardinality. We similarly define maximum minimal generating set having greatest possible cardinality.
The following is a generalization of Theorem 2.11 in [10] for number fields.
Theorem 7.6.
Let $K$ be a number field and $n>2$ an odd composite number. Then there exists an $L/K$ of degree $n$ for which the Galois closure has minimal generating sets of cardinalities $2,3,\dots,\omega(n)$ (where $\omega(n)$ is the number of distinct prime divisors of $n$) and these are the only possible cardinalities for minimal generating set of $\tilde{L}$.
Proof.
By Theorem 2.11 in [10] we have the result for $K=\mathbb{Q}$ that is we have $L^{\prime}/\mathbb{Q}$ satisfying the required properties. Now Galois group of Galois closure of $L^{\prime}/\mathbb{Q}$ is $G=\mathbb{Z}/n\mathbb{Z}\rtimes(\mathbb{Z}/n\mathbb{Z})^{\times}$ which is solvable. Also $H=Gal(\tilde{L^{\prime}}/L^{\prime})=1\times(\mathbb{Z}/n\mathbb{Z})^{\times}$.
By Lemma 6.6 (2), $G$ is realizable as a Galois group over $K$ say for $E/K$. Let $L/K$ be fixed field of $H$ inside $E/K$. By Remark 7.1.1, $L/K$ satisfies the required properties. ∎
Theorem 7.7.
Let $K$ be a number field. Given positive integers $n>2$ and $s|n$ with $s<n$. There exists an $L/K$ of degree $n$ for which the Galois closure has a minimal generating set of cardinality $s$.
Furthermore that $L/K$ satisfies $s_{K}(L)=s$. Hence there is a unique minimal generating set for the Galois closure of $L/K$ which is thus, also a minimum minimal generating set.
Proof.
Let $r=n/s$. Hence $r>1$. By the proof of second approach of Theorem 5.2 we have $L/K$ such that $[L:K]=n$ and $r_{K}(L)=r$. Thus $s_{K}(L)=s$. Let $L_{i}$’s be the $s$ many distinct fields isomorphic to $L$ over $K$. Now by Remark 5.4.2 we have for every permutation $(i_{1},i_{2},\dots,i_{s})$ of $(1,2,\dots,s)$ that
$K\subseteq L_{i_{1}}\subseteq L_{i_{1}}L_{i_{2}}\subseteq\dots\subseteq L_{i_{1}}L_{i_{2}}\cdots L_{i_{s}}$
is a cluster tower for $L/K$ of length $s+1$. Thus by Lemma 7.3, $S=\{L_{i}\}_{i=1}^{s}$ is a minimal generating set of the Galois closure of $L/K$ which is the unique minimal generating set and thus also a minimum minimal generating set.∎
Remark 7.7.1.
Assuming the condition $s\neq n$ is necessary in the above theorem. Suppose $L/K$ is degree $n$ extension. Then we can’t have a minimal generating set of cardinality $n$. This follows from the argument in Remark 5.2.1 as cardinality of minimal generating set being $n$ forces $s_{K}(L)=n$ and $r_{K}(L)=1$. Any subset of minimal generating set of cardinality n1n-1 also generates the Galois closure which contradicts the minimality.
Proposition 7.8.
Consider $L/K$.
  1. (1)
    Let $B$ be any minimal generating set of the Galois closure of $L/K$. Then $B\supset\{L_{i}\ |\ L_{S\backslash L_{i}}\neq\tilde{L}\}$.
  2. (2)
    Suppose $B\subset S$ is the unique minimal generating set of the Galois closure of $L/K$. Then $B=\{L_{i}\ |\ L_{S\backslash L_{i}}\neq\tilde{L}\}$.
Proof.
  1. (1)
    Let $L_{i}$ be such that $L_{S\backslash L_{i}}\neq\tilde{L}$. Assume on the contrary that $L_{i}\not\in B$. Thus $B\subset S\backslash L_{i}$. Thus $L_{B}\subset L_{S\backslash L_{i}}$. Hence $L_{B}\neq\tilde{L}$ which is a contradiction.
  2. (2)
    Consider $L_{i}$ such that $S\backslash L_{i}=\tilde{L}$. Now since $S\backslash L_{i}$ is a generating set of $\tilde{L}$, by Lemma 7.2 we have a minimal generating set $B^{\prime}$ of $\tilde{L}$ such that $B^{\prime}\subset S\backslash L_{i}$. By uniqueness of $B$ we have $B=B^{\prime}$. Thus $B\subset S\backslash L_{i}$. Hence $L_{i}\not\in B$. By part (1) we are done.
8. Compositum Feasible Triplets
The following notions were introduced by Drungilas et al. in [6] with $\mathbb{Q}$ as the base field $K$.
Definition 8.1.
Let $K$ be a perfect field. A triplet of positive integers $(a,b,c)$ is said to be
  • compositum feasible over $K$ if there exist $L/K$ an $L^{\prime}/K$ of degrees $a$ and $b$ respectively with the compositum $LL^{\prime}/K$ having degree $c$.
  • sum feasible over $K$ if there exist $\alpha,\beta,\gamma\in\bar{K}$ with degrees of respective minimal polynomial over $K$ as $a,b$ and $c$ such that $\alpha+\beta+\gamma=0$.
  • product feasible over $K$ if there exist $\alpha,\beta$ and $\gamma\in\bar{K}$ with degrees of respective minimal polynomial over $K$ as $a,b$ and $c$ such that $\alpha\beta\gamma=1$.
Remark 8.1.1.
If $(a,b,c)$ is compositum feasible over $K$ then clearly the permutation $(b,a,c)$ is also compositum feasible over $K$. If $(a,b,c)$ is sum feasible (or product feasible) over $K$ then clearly all the six permutations are also sum feasible (or product feasible) over $K$.
The proof of Proposition 1 in [6] also works for any infinite perfect field as base field in place of $\mathbb{Q}$. We state that result for infinite perfect fields below.
Lemma 8.2.
Let $K$ be an infinite perfect field. If the triplet $(a,b,c)$ is compositum feasible over $K$, then it is also sum feasible and product feasible over $K$.
Lemma 8.3 (part of Lemma 14 [6]).
Suppose that $(a,b,c)$ is compositum feasible over $K$. Then $c=lcm(a,b)\cdot t$ for some positive integer $t\leq gcd(a,b)$.
Remark 8.3.1.
If $(a,b,c)$ is compositum feasible over $K$ and some permutation other than $(b,a,c)$ is also compositum feasible over $K$ then by Lemma 8.3, we have $c=a$ or $c=b$. If all permutations are compositum feasible over $K$ then $a=b=c$.
Proposition 8.4.
Suppose $(a,b,c)$ is compositum feasible over $K$.
  1. (1)
    Suppose $a<c$ and $b<c$. Then the Galois group of the Galois closure of the compositum field over $K$ cannot be $S_{c}$ or $A_{c}$.
  2. (2)
    Suppose the Galois group of the Galois closure of the compositum field over $K$ is cyclic. Then $c=lcm(a,b)$.
Proof.
  1. (1)
    Assume on the contrary that we have $L/K$ and $L^{\prime}/K$ of degrees $a$ and $b$ such that compositum $LL^{\prime}/K$ has degree $c$ and Galois closure of $LL^{\prime}/K$ say $M/K$ has Galois group $S_{c}$ or $A_{c}$. In that case, the subgroup fixing $LL^{\prime}$ is isomorphic to $S_{c-1}$ or $A_{c-1}$ respectively. Since this subgroup is maximal in the whole group, we cannot have nontrivial intermediate fields of $LL^{\prime}/K$, which is a contradiction.
  2. (2)
    Suppose we have $L/K$ and $L^{\prime}/K$ of degrees $a$ and $b$ such that compositum $LL^{\prime}/K$ has degree $c$ and Galois closure of $LL^{\prime}/K$ say $M/K$ has cyclic Galois group. Then clearly $M=LL^{\prime}$. Now $Gal(M/K)=\mathbb{Z}/c\mathbb{Z}=<\bar{1}>$. Thus $Gal(M/L)=\mathbb{Z}/(c/a)\mathbb{Z}=<\bar{a}>$ and $Gal(M/L^{\prime})=\mathbb{Z}/(c/b)\mathbb{Z}=<\bar{b}>$. Since $M=LL^{\prime}$, we have $\bar{a}>\cap<\bar{b}>=\{\bar{0}\}$. We have $c\geq lcm(a,b)$. Assume $c>lcm(a,b)$. Then $\bar{0}\neq\bar{lcm(a,b)}\in<\bar{a}>\cap<\bar{b}>$ which is a contradiction. Hence $c=lcm(a,b)$.
The following is Theorem 2.4 in [1] stated here for $l=2$ and base field any perfect field $K$ in place of $\mathbb{Q}$ (the same proof still works). The second condition in the hypothesis is stated as in Remark 2.4.1 (2) in [1].
Lemma 8.5.
Let $K$ be a perfect field and $J/K$ be a finite extension and let $E_{1}$ and $E_{2}$ be finite Galois extensions of $K$ contained in $\bar{K}$. Let $G_{i}=Gal(E_{i}/K)$ for $i=1,2$. Suppose
  1. (1)
    $E_{1}\cap E_{2}=J$
  2. (2)
    for every $i$, $G_{i}\cong H_{i}\rtimes G_{i}/H_{i}$ (for some semidirect product group law) where $H_{i}=Gal(E_{i}/J)$.
Then $Gal(E_{1}E_{2}/K)\cong(H_{1}\times H_{2})\rtimes G_{1}/H_{1}$ (for some semidirect product group law).
Lemma 8.6.
A triplet of positive integers $(a,b,c)$ satisfies $max\{v_{p}(a),v_{p}(b)\}\leq v_{p}(c)\leq v_{p}(a)+v_{p}(b)$ for every prime number $p$ (where $v_{p}$ is usual $p$-adic valuation) if and only if $c=lcm(a,b)\cdot t$ where t|gcd(a,b)t\ |\ gcd(a,b).
Proof.
Suppose $c=lcm(a,b)\cdot t$ where t|gcd(a,b)t\ |\ gcd(a,b). Thus $c=ab/s$ where s|gcd(a,b)s\ |\ gcd(a,b). We have $v_{p}(c)=v_{p}(a)+v_{p}(b/s)=v_{p}(a/s)+v_{p}(b)$ for every prime. Hence the triplet clearly satisfies the given condition.
Conversely suppose that the triplet satisfies the condition. Now $a=\Pi p^{v_{p}(a)},b=\Pi p^{v_{p}(b)},$
$c=\Pi p^{v_{p}(c)},lcm(a,b)=\Pi p^{max\{v_{p}(a),v_{p}(b)\}},gcd(a,b)=\Pi p^{min\{v_{p}(a),v_{p}(b)\}}$ and $ab=\Pi p^{v_{p}(a)+v_{p}(b)}$. Since $max\{v_{p}(a),v_{p}(b)\}\leq v_{p}(c)$ for every prime $p$. Hence $lcm(a,b)|c$.
Now $t=c/lcm(a,b)=\Pi p^{v_{p}(c)-max\{v_{p}(a),v_{p}(b)\}}$. Since $v_{p}(c)\leq v_{p}(a)+v_{p}(b)=max\{v_{p}(a),v_{p}(b)\}+min\{v_{p}(a),v_{p}(b)\}$ for every prime number $p$. Thus t|gcd(a,b)t\ |\ gcd(a,b).∎
Remark 8.6.1.
Clearly the statements in the above lemma are also equivalent to lcm(a,b)|c|ablcm(a,b)\ |\ c\ |\ ab.
It follows from Lemma 8.6 that the following result that we establish is a generalization of Theorem 7 in [6] for number fields.
Theorem 8.7.
Let $K$ be a number field. Given positive integers $a,b$ and t|gcd(a,b)t\ |\ gcd(a,b) we have that $(a,b,lcm(a,b)\cdot t)$ is a compositum feasible triplet over $K$. Furthermore we can choose the field extensions such that compositum has Galois closure over $K$ having a solvable Galois group.
We also have that $(a,b,lcm(a,b)\cdot t)$ is a sum feasible and product feasible triplet over $K$.
Proof.
Since t|gcd(a,b)t\ |\ gcd(a,b), we have $gcd(a,b)=st$ for some positive integer $s$. Thus $lcm(a,b)\cdot t=ab/s$. Thus the triplet is $(a,b,ab/s)$ for s|gcd(a,b)s\ |\ gcd(a,b).
If $s=a$ or $b$. Then the triplet is $(a,b,b)$ with $a|b$ or $(a,b,a)$ with $b|a$. Without loss of generality let $a\leq b$ and $a|b$. By Lemma 5.4 for $K/K$, we get a cyclic Galois extension $L^{\prime}/K$ of degree $b$. Since $a|b$ and $L^{\prime}/K$ is cyclic, we have an intermediate extension $L/K$ of degree $a$. Thus clearly $LL^{\prime}=L^{\prime}$ which is Galois over $K$ with a cyclic hence a solvable Galois group.
If $s<a,b$. Then $a,b>1$. Let $r=a/s$. Thus $r>1$. For $a=2$ we have $s=1$ and $r=2$. By Lemma 5.4 for $K/K$, we get a cyclic Galois extension $L/K$ of degree $2$ and hence $r_{K}(L)=2$ and $G_{1}=Gal(\tilde{L}/K)$ is $\mathbb{Z}/2\mathbb{Z}$. For $a>2$, we proceed as we did in the proof of Theorem 5.1 in [2]. Thus we have that there exists an extension $L/K$ with Galois closure $\tilde{L}$ such that $[L:K]=a$, $r_{K}(L)=r$ and $G_{1}=Gal(\tilde{L}/K)$ is $({\mathbb{Z}}/r{\mathbb{Z}})^{s}\rtimes{\mathbb{Z}}/s{\mathbb{Z}}$.
By Theorem 7.3.4 in [2], we get an intermediate extension $N/K$ of $L/K$ such that $L/N$ is cyclic Galois extension of degree $r$ and $N/K$ is cyclic Galois extension of degree $s$ and $H_{1}=Gal(\tilde{L}/N)$ is $({\mathbb{Z}}/r{\mathbb{Z}})^{s}$. Now $b/s>1$. Hence by Lemma 5.4 for $\tilde{L}/K$, there exists a cyclic Galois extension $F/K$ of degree $b/s$ such that $\tilde{L}$ and $F$ are linearly disjoint over $K$.
Let $L^{\prime}=NF$. Since $\tilde{L}\cap F=K$. Thus $N\cap F=K$. Hence $L^{\prime}:N]=[F:K]=b/s$ and $L^{\prime}/K$ is Galois with $L^{\prime}:K]=[L^{\prime}:N][N:K]=(b/s)\cdot s=b$ and $G_{2}=Gal(L^{\prime}/K)=(\mathbb{Z}/(b/s)\mathbb{Z})\times(\mathbb{Z}/s\mathbb{Z})$ and $H_{2}=Gal(L^{\prime}/N)=\mathbb{Z}/(b/s)\mathbb{Z}$. We also have $\tilde{L}\cap L^{\prime}=N$. Hence $LL^{\prime}:N]=[L:N][L^{\prime}:N]=r\cdot b/s=ab/s^{2}$. Thus $LL^{\prime}:K]=[LL^{\prime}:N][N:K]=(ab/s^{2})\cdot s=ab/s$.
By Lemma 8.5 for $J=N$, $E_{1}=\tilde{L}$ and $E_{2}=L^{\prime}$, we get that $Gal(\tilde{L}L^{\prime}/K)$ is $(H_{1}\times H_{2})\rtimes G_{1}/H_{1}=(({\mathbb{Z}}/(a/s){\mathbb{Z}})^{s}\times\mathbb{Z}/(b/s)\mathbb{Z})\rtimes{\mathbb{Z}}/s{\mathbb{Z}}$ (for some semidirect product group law). This group is clearly solvable. Since $\tilde{L}L^{\prime}$ is Galois closure of $LL^{\prime}/K$, we are done.
The last assertion follows from Lemma 8.2.∎
Remark 8.7.1.
Theorem 8.7 doesn’t cover the case $(a,b,lcm(a,b)\cdot t)$ where $t\leq gcd(a,b)$ and $t\nmid gcd(a,b)$. In fact it is not necessary that such a triplet is compositum feasible over number fields. For example (p,p,p(pl))(p,p,p(p-l)) is not compositum feasible over $\mathbb{Q}$ for every integer $l\geq 2$ and every prime $p>l^{2}-l+1$ (See Theorem 1.2 [5]). But we do have examples of such triplets like (n,n,n(n1))(n,n,n(n-1)) for $n\geq 3$ which are indeed compositum feasible over $\mathbb{Q}$ (See Proposition 29 (1) [6]).
Remark 8.7.2.
Given any compositum feasible triplet $(a,b,c)$ over a perfect field $K$, it is is not always necessary that we can choose the field extensions such that compositum has Galois closure over $K$ having a solvable Galois group.
As noted in [17] (See note after Lemmas 2.7 and 2.8 in [17]), we have for $p$ prime and $1\leq s\leq p-1$ that if $(p,p,ps)$ is compositum feasible over $\mathbb{Q}$ for field extensions with Galois group of Galois closure of compositum field as solvable then we should have s|(p1)s\ |\ (p-1).
We have from Theorem 1.1 (c) in [17] that $(11,11,11\cdot 6)$ is compositum feasible over $\mathbb{Q}$. Since $6\nmid 10$. Hence we cannot choose the field extensions for the triplet $(11,11,11\cdot 6)$ such that the Galois closure of the compositum has a solvable Galois group.
The following cases (some of which may be covered over $\mathbb{Q}$ in earlier works on this problem) follow easily (over any number field) from Theorem 8.7.
Corollary 8.7.1.
We have the following triplets to be compositum feasible over any number field.
  1. (1)
    $(n,n,nd)$ for positive integers $n$ and $d|n$. In particular $(p^{k},p^{k},p^{j})$ for positive integers $p,k$ and $j$ where $p$ is prime and $k\leq j\leq 2k$.
  2. (2)
    $(p,t,t)$ for positive integers $p$ and $t$ where $p$ is prime and $p|t$ ([6]).
  3. (3)
    $(a(abc)^{n},b(abc)^{n},c(abc)^{n+1})$ for positive integers $a,b$ and $c$ and $n\geq 2$ ([6]).
The following follows from Lemma 6.5.
Proposition 8.8.
Suppose $(a,b,c)$ is compositum feasible over $\mathbb{Q}$. If we can choose the field extensions such that compositum has Galois closure over $\mathbb{Q}$ having Galois group $G$ such that for every positive integer $n$, the $n$-fold direct product $G^{n}$ occurs as a Galois group over $\mathbb{Q}$. Then $(a,b,c)$ is compositum feasible over any number field.
Corollary 8.8.1.
Suppose $(a,b,c)$ is compositum feasible over $\mathbb{Q}$. If we can choose the field extensions such that compositum has Galois closure over $\mathbb{Q}$ having a solvable Galois group. Then $(a,b,c)$ is compositum feasible over any number field.
We establish Proposition 8.10 which is a partial case of Conjecture 4 in [6] where it is conjectured that the set of compositum feasible triplets over $K$, which we denote by $\mathcal{C}_{K}$, is a commutative multiplicative monoid with $(1,1,1)$ as the identity element.
Remark 8.8.1.
Clearly if $(a,b,c),(a^{\prime},b^{\prime},c^{\prime})\in\mathcal{C}_{K}$ with $gcd(c,c^{\prime})=1$. Then $(aa^{\prime},bb^{\prime},cc^{\prime})\in\mathcal{C}_{K}$. Suppose $L/K$ and $L^{\prime}/K$ have degrees $a$ and $b$ such that $LL^{\prime}/K$ has degree $c$ and $M/K$ and $M^{\prime}/K$ have degrees $a^{\prime}$ and $b^{\prime}$ such that $MM^{\prime}/K$ is of degree $c^{\prime}$. Since $gcd(c,c^{\prime})=1$ and lcm(a,b)|clcm(a,b)\ |\ c and $lcm(a^{\prime},b^{\prime})\ |\ c^{\prime}$. Thus $gcd(a,a^{\prime})=gcd(b,b^{\prime})=1$. Thus degrees of $LM/K$, $L^{\prime}M^{\prime}/K$ and $(LL^{\prime})(L^{\prime}M^{\prime})/K$ are $aa^{\prime},bb^{\prime}$ and $cc^{\prime}$ respectively and $(LL^{\prime})(MM^{\prime})=(LM)(L^{\prime}M^{\prime})$.
Lemma 8.9 (Proposition 3.2 [4]).
Suppose $(a,b,c),(a^{\prime},b^{\prime},c^{\prime})\in\mathcal{C}_{K}$. If for one of the triplets we can choose the field extensions such that compositum has Galois closure over $K$ having Galois group $G$ such that for every positive integer $n$, the $n$-fold direct product $G^{n}$ occurs as a Galois group over $K$. Then $(aa^{\prime},bb^{\prime},cc^{\prime})\in\mathcal{C}_{K}$.
Corollary 8.9.1.
Let $K$ be a number field. Suppose $(a,b,c),(a^{\prime},b^{\prime},c^{\prime})\in\mathcal{C}_{K}$. If for one of the triplets we can choose the field extensions such that compositum has Galois closure over $K$ having a solvable Galois group. Then $(aa^{\prime},bb^{\prime},cc^{\prime})\in\mathcal{C}_{K}$.
Let $\mathcal{C}^{\prime}_{K}$ be the set consisting of triplets of the form $(a,b,c)$ with $c=lcm(a,b)\cdot t$ where t|gcd(a,b)t\ |\ gcd(a,b).
Proposition 8.10.
Let $K$ be a number field.
  1. (1)
    Suppose $(a,b,c),(a^{\prime},b^{\prime},c^{\prime})\in\mathcal{C}_{K}$. If $(a,b,c)\in\mathcal{C}^{\prime}_{K}$, then we have that $(aa^{\prime},bb^{\prime},cc^{\prime})\in\mathcal{C}_{K}$.
  2. (2)
    We have that $\mathcal{C}^{\prime}_{K}\subset\mathcal{C}_{K}$ and $\mathcal{C}^{\prime}_{K}$ is a commutative multiplicative monoid. In particular if $(a,b,c)\in\mathcal{C}^{\prime}_{K}$ then for any $n\in\mathbb{N}$ we have $(a^{n},b^{n},c^{n})\in\mathcal{C}^{\prime}_{K}$.
Proof.
  1. (1)
    Follows from Theorem 8.7 and Corollary 8.9.1.
  2. (2)
    Suppose $(a,b,c),(a^{\prime},b^{\prime},c^{\prime})\in\mathcal{C}^{\prime}_{K}$. Thus $c=ab/s$ where s|gcd(a,b)s\ |\ gcd(a,b) and $c^{\prime}=a^{\prime}b^{\prime}/s^{\prime}$ where $s^{\prime}\ |\ gcd(a^{\prime},b^{\prime})$ as in proof of Theorem 8.7. Consider $(aa^{\prime},bb^{\prime},cc^{\prime})$. Now $cc^{\prime}=(aa^{\prime})(bb^{\prime})/(ss^{\prime})$. Since we have $ss^{\prime}\ |\ gcd(a,b)gcd(a^{\prime},b^{\prime})$ and $gcd(a,b)gcd(a^{\prime},b^{\prime})\ |\ gcd(aa^{\prime},bb^{\prime})$, we are done.
Lemma 8.11 (Theorem 1.1 [4]).
If the triplet $(a,b,c)$ is sum feasible over $\mathbb{Q}$, then it is also product feasible over $\mathbb{Q}$.
Proposition 8.12.
Let $K$ be a number field. Let $n\in\mathbb{N}$ and $i,j$ and $k$ be non-negative integers.
  1. (1)
    Suppose $i\leq j\leq k$ and $k+i\leq n-1$. Then $(^{n}P_{j},^{n}P_{k},^{n}P_{k+i})\in\mathcal{C}_{K}$. We have $(^{n}P_{j},^{n}P_{k},^{n}P_{k+i})\in\mathcal{C}^{\prime}_{K}$ if and only if $^{n-k}P_{i}\ |\ ^{n}P_{j}$. In particular, for $n\geq 3$ we have $(n,n,n(n-1))\in\mathcal{C}_{K}\backslash\mathcal{C}^{\prime}_{K}$.
  2. (2)
    $(^{n}C_{j},^{n}C_{k},^{n}C_{k+j})$ is sum feasible over $K$ for k+j<n1k+j<n-1. It is also product feasible over $\mathbb{Q}$. In particular, for $n>3$ we have that (n,n,n(n1)/2)(n,n,n(n-1)/2) is sum feasible over $K$.
Proof.
Let $f$ over $K$ be irreducible of degree $n>2$ with Galois group ${\mathfrak{S}}_{n}$ (Such a polynomial exists by Lemma 5.3).
  1. (1)
    Consider $k+i$ many distinct roots of $f$ in a fixed order. Let $L$ be field generated by first $j$ many roots and $L^{\prime}$ be field generated by last $k$ many roots. Then $LL^{\prime}$ is field generated by all $k+i$ many roots. We know that degree of the field generated by $m$ many distinct roots over $K$ is $^{n}P_{m}$. Thus $(^{n}P_{j},^{n}P_{k},^{n}P_{k+i})\in\mathcal{C}_{K}$. Now $(^{n}P_{j},^{n}P_{k},^{n}P_{k+i})\in\mathcal{C}^{\prime}_{K}$ $\iff$ $(^{n}P_{k+i}/^{n}P_{k})=\ ^{n-k}P_{i}$ divides $^{n}P_{j}$.
  2. (2)
    Consider $k+j$ many distinct roots of $f$ in a fixed order. Let $\alpha$ be sum of first $j$ many roots and $\beta$ be sum of last $k$ many roots. Let $\gamma$ be sum of all $k+j$ many roots. We know by proof of Theorem 7.3.6 in [2] that degree of the field generated by sum of m<n1m<n-1 many distinct roots over $K$ is $^{n}C_{m}$. Thus $(^{n}C_{j},^{n}C_{k},^{n}C_{k+j})$ is sum feasible over $K$. By Lemma 8.11 product feasibility over $\mathbb{Q}$ follows.
Remark 8.12.1.
Let $K$ be a number field. If $(a,b,c)\in\mathcal{C}^{\prime}_{K}$ and $(a^{\prime},b^{\prime},c^{\prime})\in\mathcal{C}_{K}\backslash\mathcal{C}^{\prime}_{K}$ then its neither necessary that $(aa^{\prime},bb^{\prime},cc^{\prime})$ lies in $\mathcal{C}^{\prime}_{K}$ nor necessary that it lies in $\mathcal{C}_{K}\backslash\mathcal{C}^{\prime}_{K}$. Let $n\geq 3$. Now we have $(n,n,n(n-1))\in\mathcal{C}_{K}\backslash\mathcal{C}^{\prime}_{K}$ and $(n,n,n),(n-1,n-1,n-1)\in\mathcal{C}^{\prime}_{K}$. Observe that $(n^{2},n^{2},n^{2}(n-1))\in\mathcal{C}_{K}\backslash\mathcal{C}^{\prime}_{K}$ but on the other hand $(n(n-1),n(n-1),n(n-1)^{2})\in\mathcal{C}^{\prime}_{K}$.
In fact if $(a,b,c)\in\mathcal{C}_{K}\backslash\mathcal{C}^{\prime}_{K}$ then there exist infinitely many elements of $\mathcal{C}^{\prime}_{K}$ namely $(cn,cn,cn)$ for $n\in\mathbb{N}$ such that $(a\cdot cn,b\cdot cn,c\cdot cn)\in\mathcal{C}^{\prime}_{K}$. Similarly there exist infinitely many elements of $\mathcal{C}^{\prime}_{K}$ namely $(p,p,p)$ for $p>c$ prime such that $(a\cdot p,b\cdot p,c\cdot p)\in\mathcal{C}_{K}\backslash\mathcal{C}^{\prime}_{K}$.
Similarly if $(a,b,c),(a^{\prime},b^{\prime},c^{\prime})\in\mathcal{C}_{K}\backslash\mathcal{C}^{\prime}_{K}$ such that $(aa^{\prime},bb^{\prime},cc^{\prime})\in\mathcal{C}_{K}$. Then its neither necessary that $(aa^{\prime},bb^{\prime},cc^{\prime})$ lies in $\mathcal{C}^{\prime}_{K}$ nor necessary that it lies in $\mathcal{C}_{K}\backslash\mathcal{C}^{\prime}_{K}$. For $K=\mathbb{Q}$, we have $(3,3,6),(4,4,12)\in\mathcal{C}_{K}\backslash\mathcal{C}^{\prime}_{K}$. Clearly $(12,12,72)\in\mathcal{C}^{\prime}_{K}$. By Theorem 1 in [7], we have that $(9,9,36)\in\mathcal{C}_{K}$ but clearly $(9,9,36)\not\in\mathcal{C}^{\prime}_{K}$.
The following notion was introduced in Appendix A in [13].
Definition 8.13.
[…]
Proposition 8.14.
Let $K$ be a number field and $(a,b,c)\in\mathcal{C}^{\prime}_{K}$ that is $c=lcm(a,b)\cdot t$ where t|gcd(a,b)t\ |\ gcd(a,b). Then $(a,b,c)$ is irreducible if and only if it is one of the following where $p$ is any prime:
(i) $(1,1,1)$ (ii) $(1,p,p)$ (iii) $(p,1,p)$ (iv) $(p,p,p)$.
Proof.
We will prove the result step by step. Suppose $(a,b,c)$ is irreducible.
  1. (1)
    Then we have $t=1$.
[…]
  2. (2)
    Suppose $a,b>1$, then for any prime $p$, we have $p|a$ if and only if $p|b$.
[…]
  3. (3)
    Suppose $a,b>1$, then $a$ is a prime power.
[…]
  4. (4)
    Suppose $a,b>1$, then $(a,b,c)=(p,p,p)$ for some prime $p$.
    By parts (2) and (3) we have $(a,b,c)=(p^{i},p^{j},p^{max\{i,j\}})$. Without loss of generality let $i\leq j$. Then $(a,b,c)=(p^{i}\cdot 1,p^{i}\cdot p^{j-i},p^{i}\cdot p^{j-i})$. Thus $i=0$ or $j=i$. If $i=0$ then by part (2) we have $j=0$ which is a contradiction. If $j=i$ then $(a,b,c)=(p^{i},p^{i},p^{i})$. We must have $i=1$ because if $i>1$ then $(a,b,c)=(p\cdot p^{i-1},p\cdot p^{i-1},p\cdot p^{i-1})$ which is a contradiction.
  5. (5)
    Finally if $a=1$ or $b=1$ then $(a,b,c)=(1,1,1)$ or $(1,p,p)$ or $(p,1,p)$ for some prime $p$.
    It is enough to show for $a=1$ and $b>1$. If $b$ is not prime then $b=de$ where $d,e>1$. Then $(a,b,c)=(1\cdot 1,d\cdot e,d\cdot e)$ which is a contradiction.
Remark 8.14.1.
The above proof also demonstrates that whenever $(a,b,c)\in\mathcal{C}^{\prime}_{K}$ is reducible inside $\mathcal{C}_{K}$, it is also reducible inside $\mathcal{C}^{\prime}_{K}$ (product of two triplets in $\mathcal{C}^{\prime}_{K}$).
Remark 8.14.2.
If we do not have the condition that t|gcd(a,b)t\ |\ gcd(a,b), then it is not necessary that $(a,b,c)$ being irreducible implies $t=1$. For instance the triplet (n,n,n(n1))(n,n,n(n-1)) for $n\geq 3$ which is compositum feasible over $\mathbb{Q}$ is irreducible as shown in Proposition A1 in [13].
Proposition 8.15.
Let $K$ be a number field. Then
  1. (1)
    For any $n\in\mathbb{N}$, $(n,n,n\phi(n))\in\mathcal{C}_{K}$ where $\phi$ is the Euler totient function. Furthermore we can choose the field extensions such that compositum has Galois closure over $K$ having a solvable Galois group.
  2. (2)
    $(n,n,n\phi(n))\in\mathcal{C}^{\prime}_{K}$ if and only if $n=1$ or $n=2^{l}3^{m}$ where $l\geq 1$ and $m\geq 0$.
  3. (3)
    $(n,n,n\phi(n))$ is irreducible if and only if $n=1$ or $n$ is a prime.
Proof.
  1. (1)
    First let $K=\mathbb{Q}$. Consider the case in Example 2.3 (1). Let $L=\mathbb{Q}(a)$ and $L^{\prime}=\mathbb{Q}(a\zeta)$. Thus $LL^{\prime}=\mathbb{Q}(a,\zeta)=\tilde{L}$. Since the Galois group of Galois closure over $\mathbb{Q}$ is solvable, we are done for any number field $K$ by Lemma 6.6 (2).
  2. (2)
    Now $(n,n,n\phi(n))\in\mathcal{C}^{\prime}_{K}$ if and only if $\phi(n)\ |\ n$ that is $\Pi_{p|n}(p-1)\ |\ \Pi_{p|n}(p)$. It is easy to see that this holds if and only if $n=1$ or $n=2^{l}3^{m}$ for $l\geq 1$ and $m\geq 0$.
  3. (3)
    Suppose the triplet is irreducible and $n\neq 1$. Let $p|n$ and $n=p^{k}m$ where $p$ is a prime and $p\nmid m$. Therefore $(n,n,n\phi(n))=(p^{k}m,p^{k}m,p^{k}m\ \phi(p^{k}m))=(p^{k}\cdot m,p^{k}\cdot m,p^{k}\phi(p^{k})\cdot m\phi(m))$. Thus we must have $m=1$. Now $n=p^{k}$. Thus $(n,n,n\phi(n))=(p^{k},p^{k},p^{k}(p^{k}-p^{k-1}))=(p\cdot p^{k-1},p\cdot p^{k-1},p^{2}\cdot p^{k-1}(p^{k-1}-p^{k-2}))$. Hence we must have $k=1$. The converse is clear.
Remark 8.15.1.
In particular Proposition 8.15 says that for $p\geq 3$ prime we have $(p,p,p(p-1))\in\mathcal{C}_{K}$ and we can choose the field extensions such that compositum has Galois closure over $K$ having a solvable Galois group. Also by Proposition 8.12 we have that for the same triplet $(p,p,p(p-1))\in\mathcal{C}_{K}$ we can choose the field extensions such that compositum has Galois closure over $K$ having Galois group $\mathfrak{S}_{p}$ which is non-solvable for $p\geq 5$.
9. Improving on the Inverse Problems
The following lemma follows from the proof in [2] of Lemma 6.5 (essentially the proof of Theorem 4.2 in [3]). We provide the proof for the sake of completeness.
Lemma 9.1.
Suppose for a group $G$, direct product $G^{m}$ is realizable as a Galois group over $K$ for an $m\in\mathbb{N}$. Then we get $m$-many Galois extensions of $K$ inside $\bar{K}$ which are pairwise non-isomorphic over $K$ and are pairwise linearly disjoint over $K$ with each having Galois group $G$ over $K$.
Proof.
Now $G^{m}$ is realizable over $K$, say for $E/K$ Galois, we have ${\rm Gal}(E/K)\cong G^{m}$. We have normal subgroups $N_{i}=G\times G\times\dots\times{1}\times\dots\times G$ of $G^{m}$ for $1\leq i\leq m$ where the $i$th coordinate is trivial and there is no restriction in other coordinates. So $N_{i}\cong G^{m-1}$. Let $E_{i}$ be the subfield of $E$ corresponding to $N_{i}$, so $E_{i}/K$ is Galois with ${\rm Gal}(E_{i}/K)\cong G^{m}/N_{i}\cong G$.
We observe that $N_{i}$ are not conjugate to each other in $G^{m}$ and they pairwise generate $G^{m}$. Hence $E_{i}$ are not isomorphic to each other over $K$ and are pairwise linearly disjoint over $K$. ∎
The following is an improvement on Theorem 5.1.
Theorem 9.2.
Let $K$ be a number field. Let $n>2$ and $r|n$. Then we get arbitrarily large finite families of extensions $L/K$ inside $\bar{K}$ which are pairwise non-isomorphic over $K$ and are pairwise linearly disjoint over $K$ and each having degree $n$ with cluster size $r_{K}(L)=r$.
Proof.
We proceed in the same way as the author and Bhagwat proceeded in the proof of Theorem 5.1 in [2]. Suppose $r=1$. By Lemma 5.3, there exist infinitely many pairwise linearly disjoint extensions over $K$ of degree $n$ with Galois closures having Galois group ${\mathfrak{S}}_{n}$ over $K$. Hence we obtain infinitely many extensions of degree $n$ which are pairwise non-isomorphic and pairwise linearly disjoint over $K$ and these extensions have cluster size 1.
Suppose $r>1$. We have that there exists $L/K$ with $G=Gal(\tilde{L}/K)$ solvable and $[L:K]=n$ and $r_{K}(L)=r$. Consider subgroup $H=Gal(\tilde{L}/L)\subset G$. Since direct product of solvable groups is solvable, direct products $G^{m}$ for $m\in\mathbb{N}$ are solvable. By Shafarevich’s theorem ([16]), $G^{m}$ for $m\in\mathbb{N}$ are realizable as Galois groups over $\mathbb{Q}$. Hence by Lemma 6.5, $G^{m}$ for $m\in\mathbb{N}$ are realizable as Galois groups over number field $K$.
By Lemma 9.1, for any $m\in\mathbb{N}$, we get $E_{i}$ for $1\leq i\leq m$ which are pairwise non-isomorphic over $K$ and are pairwise linearly disjoint over $K$ with each having Galois group $G$ over $K$. Thus for $1\leq i\leq m$, we have $E_{i}^{H}$ (which correspond to subgroups $G\times G\times\dots\times H\times\dots\times G$ of $G^{m}$ where the $i$th coordinate is an element of $H$ and there is no restriction in other coordinates) which are pairwise non-isomorphic over $K$ and are pairwise linearly disjoint over $K$ with each having Galois closures with Galois group $G$ over $K$. Each $E_{i}^{H}/K$ has degree $n$ and cluster size rr.∎
Remark 9.2.1.
In the above proof, for any given $m\in\mathbb{N}$, we get $m$ many extensions over $K$, with degree $n$ and cluster size $r$, which are pairwise non-isomorphic over $K$. Thus for any given $m$, we get $mn/r$ many distinct extensions over $K$, with degree $n$ and cluster size $r$.
Remark 9.2.2.
The cases $r=1$ for $n$ odd and $r=2$ for $n$ even have other obvious isomorphism classes too distinct from the ones obtained in above theorem. Namely the ones considered in Example 2.3 (1) and Theorem 7.6 where Galois group of Galois closure of extension is $\mathbb{Z}/n\mathbb{Z}\rtimes(\mathbb{Z}/n\mathbb{Z})^{\times}$. Observe that even this group is solvable, so a similar story unfolds.
Similarly we have an improvement on Theorem 6.1.
Theorem 9.3.
Let $K$ be a number field. Let $n>2$ and $t|n$. Then we get arbitrarily large finite families of extensions $L/K$ inside $\bar{K}$ which are pairwise non-isomorphic over $K$ and are pairwise linearly disjoint over $K$ and each having degree $n$ with ascending index $t_{K}(L)=t$.
An improvement on Inverse root capacity problem for number fields Theorem 5.2.
Theorem 9.4.
Let $K$ be a number field. Given $(n,r,\rho)$ where $n>2$ and $r|n$ and $r|\rho$ and $\rho\neq n-1$. We get arbitrarily large finite families of degree $n$ extensions $L/K$ inside $\bar{K}$ with cluster size $r$, which are pairwise non-isomorphic over $K$ and are pairwise linearly disjoint over $K$, for which we have extensions $M/K$ such that $\rho_{K}(M,L)=\rho$.
For $\rho\neq 0$, we get $M/K$ as an extension of $L/K$ contained in $\tilde{L}$. Thus extensions $M/K$ corresponding to extensions $L/K$ which are pairwise non-isomorphic over $K$ and pairwise linearly disjoint over $K$ are themselves pairwise non-isomorphic over $K$ and pairwise linearly disjoint over $K$.
An improvement on Theorem 7.6.
Theorem 9.5.
Let $K$ be a number field and $n>2$ an odd composite number. Then we get arbitrarily large finite families of degree $n$ extensions $L/K$ inside $\bar{K}$ which are pairwise non-isomorphic over $K$ and are pairwise linearly disjoint over $K$, for which the Galois closure has minimal generating sets of cardinalities $2,3,\dots,\omega(n)$ and these are the only possible cardinalities for minimal generating set of Galois closure of $L/K$.
An improvement on Theorem 7.7.
Theorem 9.6.
Let $K$ be a number field. Given positive integers $n>2$ and $s|n$ with $s<n$. We get arbitrarily large finite families of degree $n$ extensions $L/K$ inside $\bar{K}$ which are pairwise non-isomorphic over $K$ and are pairwise linearly disjoint over $K$, for which the Galois closure has a minimal generating set of cardinality $s$.
Furthermore each $L/K$ satisfies $s_{K}(L)=s$. Hence there is a unique minimal generating set for the Galois closure of $L/K$ which is thus, also a minimum minimal generating set.
An improvement on Theorem 8.7.
Theorem 9.7.
Let $K$ be a number field. Given positive integers $a,b$ and t|gcd(a,b)t\ |\ gcd(a,b) we have that $(a,b,lcm(a,b)\cdot t)$ is a compositum feasible triplet over $K$ and we get arbitrarily large finite families of extensions $M/K$ inside $\bar{K}$ which are pairwise non-isomorphic over $K$ and are pairwise linearly disjoint over $K$ and each having degree $lcm(a,b)\cdot t$ and each is compositum of two intermediate extensions of degrees $a$ and $b$.
Acknowledgements: The author would like to thank Prof Purusottam Rath, CMI Chennai for suggesting to consider the question of isomorphism classes of extensions with given degree and cluster size which is dealt with in Section 9. The author is grateful to Prof P Vanchinathan and Dr Anand Chitrao for suggesting to generalize the main result in [10] for number fields which is achieved in a result in Section 7. The author would also like to thank Dr Shripad Garge, IIT Bombay for introducing the paper by Drungilas et al. [6] and suggesting to investigate the problem of compositum feasible triplets and associated Galois groups which is dealt with in Section 8. Finally the author would like to acknowledge support of IIT Bombay Institute Post Doctoral Fellowship during the later part of this work.